### v1

# Probabilistic Robustness Analysis in High Dimensional Space: Application to Semantic Segmentation Network

Abstract
Semantic segmentation networks (SSNs) play a critical role in domains such as medical imaging, autonomous driving, and environmental monitoring, where safety hinges on reliable model behavior under uncertainty. Yet, existing probabilistic verification approaches struggle to scale with the complexity and dimensionality of modern segmentation tasks, often yielding guarantees that are too conservative to be practical. We introduce a probabilistic verification framework that is both architecture-agnostic and scalable to high-dimensional outputs. Our approach combines sampling-based reachability analysis with conformal inference (CI) to deliver provable guarantees while avoiding the excessive conservatism of prior methods. To counteract CI’s limitations in high-dimensional settings, we propose novel strategies that reduce conservatism without compromising rigor. Empirical evaluation on large-scale segmentation models across CamVid, OCTA-500, Lung Segmentation, and Cityscapes demonstrates that our method provides reliable safety guarantees while substantially tightening bounds compared to SOTA. We also provide a toolbox implementing this technique, available on github. [^1] https://github.com/Navidhashemicodes/SSN_Reach_CLP_Surrogate
Keywords: Robustness, Semantic Segmentation, Verification, Conformal Inference
1 Introduction
Deep Neural Networks have achieved outstanding success across diverse domains, including image classification, speech recognition, medical diagnosis, and autonomous systems. However, despite their effectiveness, DNNs remain highly sensitive to input perturbations and adversarial examples—carefully crafted modifications that can cause significant changes in model predictions. This vulnerability poses serious risks, particularly in safety-critical applications such as autonomous driving, and medical imaging, where even a single misprediction may lead to catastrophic consequences.
Research on neural network robustness has generally followed two main directions. Falsification-based approaches aim to uncover adversarial examples that expose a model’s vulnerabilities (Madry et al., 2017; Goodfellow et al., 2014; Kurakin et al., 2016). While effective for identifying potential failure cases, these methods provide no formal guarantees of safety. In contrast, verification-based approaches seek to certify robustness by proving that a network remains reliable under specified input perturbations (Zhou et al., 2024; Lemesle et al., 2024; Wu et al., 2024). Although these techniques offer provable guarantees, they are typically hindered by severe scalability challenges.
Reachability analysis—the task of computing the set of all possible outputs of a system given a bounded input set—has become a central tool for verifying robustness. In the context of neural networks, it captures how input perturbations propagate through the layers and affect the final prediction. However, reachability analysis quickly becomes computationally infeasible for large-scale models or high-dimensional outputs. Semantic segmentation networks (SSNs), which assign a class label to every pixel, present a particularly difficult case due to their extremely high-dimensional output space and complex architectures, often involving components such as transposed convolutions, max-pooling, and dilated convolutions. For such models, traditional reachability-based methods either fail to scale or yield overly conservative guarantees.
To overcome these challenges, probabilistic verification has emerged as a promising direction (Fischer et al., 2021; Anani et al., 2024; Hao et al., 2022; Marzari et al., 2024). Instead of requiring guarantees to hold for every possible perturbation, this approach establishes them with provable probability under a prescribed distribution of input uncertainty. Such methods tend to scale more effectively and yield informative statistical assurances, particularly when exact reachability bounds or Lipschitz estimates are impractical or overly pessimistic. Yet, current probabilistic techniques remain limited: they often cannot capture complex output specifications and struggle to handle the scale of modern semantic segmentation networks. In this paper, we introduce a scalable, architecture-independent probabilistic verification framework with reduced conservatism, specifically designed for semantic segmentation models. Our approach makes no restrictive assumptions on network layers or structure, and is naturally applicable to high-dimensional outputs.
Related Works.
[…]
Deterministic Verification of Neural Networks. Deterministic verification of neural networks aims to provide formal guarantees that the network’s outputs meet a given specification for all inputs within a prescribed set. This problem has been addressed using several well-established techniques, such as Satisfiability Modulo Theories (SMT) solvers (Duong et al., 2023; Katz et al., 2017; Wu et al., 2024), Mixed Integer Linear Programming (MILP) formulations (Anderson et al., 2020; Cheng et al., 2017; Tjeng et al., 2017), and different types of reachability analysis (Bak et al., 2020; Tran et al., 2020). Many of these methods can be viewed through the general framework of branch-and-bound algorithms (Bunel et al., 2020), which iteratively partition and prune the input domain to either uncover counterexamples or prove correctness. Prominent tools in this area, including $\alpha,\beta$-CROWN (Zhou et al., 2024), PyRAT (Lemesle et al., 2024), and Marabou (Wu et al., 2024), have demonstrated strong results in recent VNN-COMP competitions Brix et al. (2024), consistently ranking at the top of evaluation benchmarks. A key component of branch-and-bound approaches is computing tight output bounds for neural networks over constrained input regions. Techniques like interval arithmetic (Cheng et al., 2017; Pulina and Tacchella, 2010) are commonly employed to compute these bounds, influencing both the accuracy and computational efficiency of the verification.
2 Preliminaries
Notations. We write $x\sim\mathcal{X}$ to indicate that the random variable $x$ is drawn from the distribution $\mathcal{X}$, and $x\stackrel{{\scriptstyle\mathcal{X}}}{{\sim}}\mathbf{X}$ to indicate that $x$ is sampled from the set $\mathbf{X}$ according to the distribution $\mathcal{X}$. We use bold letters to represent sets and calligraphic letters to denote distributions. Consider a 3-dimensional tensor $x\in\mathbb{R}^{h\times w\times nc}$, which we flatten into a vector. The vectorized form is denoted as $\mathbf{vec}(x)\in\mathbb{R}^{n_{0}}$, where $n_{0}=nc\times w\times h$. The Minkowski sum is denoted by $\oplus$. The ceiling operator $\lceil x\rceil$ denotes the smallest integer greater than or equal to $x\in\mathbb{R}$. A feedforward neural network with input dimension $n_{0}$, $l$ hidden layers of sizes $n_{1},\ldots,n_{l}$, and output dimension $N$, using the $\mathrm{ReLU}$ activation function, is denoted as $n_{0},n_{1},\ldots,n_{l},N](\mathrm{ReLU})$. Her we denote, a $\ell_{2}$ ball with radius $r\in\mathbb{R}^{n_{0}}$ and center $x\in\mathbb{R}^{n_{0}}$ with $\mathbf{B}_{r}(x)$.
Semantic Segmentation Neural Networks
A semantic segmentation neural network (SSN) is a nonlinear function that assigns a class label to each pixel $x(i,j)$ in a multichannel input image $x$, producing a target class label $\mathsf{mask}(i,j)$ from a predefined set of classes $\mathbf{L}=\left\{1,2,\ldots,L\right\}$,
$\mathsf{SSN}:x\in\mathbb{R}^{h\times w\times nc}\mapsto\mathsf{mask}\in\mathbf{L}^{h\times w},$
where $h$, $w$, and $nc$ denote the height, width, and number of channels of the input image, respectively, and $(i,j)\in{1,\ldots,h}\times{1,\ldots,w}$ are pixel coordinates. In this paper, we denote the logits of SSNs by $y\in\mathbb{R}^{h\times w\times L}$, where the target class label $\mathsf{mask}(i,j)$ is given by $\mathsf{mask}(i,j)=\underset{l\in\mathbf{L}}{\arg\max}\ \ y(i,j,l)$. Here, we also define the function $f:\mathbb{R}^{n_{0}}\mapsto\mathbb{R}^{n}$ as a nonlinear map between the flattened images and the flattened logits. Formally,
[…]
Conformal Inference
Consider a collection of i.i.d and positive scalar random variables $\mathbf{M}=\left\{R_{1},R_{2},\ldots,R_{m}\right\}$ sampled from $R\sim\mathcal{D}$, where $R_{1}<R_{2}<\ldots<R_{m}$. Given a new draw $R_{m+1}$ from the same distribution and a miscoverage level $\epsilon\in(0,1)$, conformal inference (CI)(Vovk et al., 2005) constructs a prediction interval $C(R_{m+1})=[0,d]$ such that $\Pr[\ \ R_{m+1}\in C(R_{m+1})\ \ ]\ \geq 1-\epsilon$. In this paper, we refer to these scalar random variables as nonconformity scores and we refer to $\mathbf{M}$ as the calibration dataset. To compute the threshold $d$, we use the empirical distribution of nonconformity scores $R_{i}$ for $i\in 1,2,\ldots,m$, and define a rank $\ell:=\lceil(m+1)(1-\epsilon)\rceil$. A valid choice for the threshold $d$ is then $R_{\ell}$, the $\ell$-th smallest element of the augmented set $R_{1},\ldots,R_{m},\infty$}, as defined in (Tibshirani et al., 2019), which yields the following marginal guarantee:
$\Pr\left[R_{m+1}\leq R_{\ell}\right]\geq 1-\epsilon,$ (1)
where the probability is over the joint randomness of both the calibration and test samples (Tibshirani et al., 2019; Vovk et al., 2005). The test sample $R_{m+1}$ typically refers to an unseen example drawn from the same distribution for which we aim to provide a coverage guarantee. In this paper, we also refer to this sample as the unseen, and denote it by $R^{\text{unseen}}$.
Given a sampled calibration dataset, the coverage level $\delta=\Pr[R^{\text{unseen}}\leq R_{\ell}]$ is itself a random variable following a $\mathrm{Beta}(\ell,m+1-\ell)$ distribution (Angelopoulos and Bates, 2021). The following equations shows the mean value and the variance of the beta distribution in terms of $m,\ell$:
$\mathbf{E}\big{[}\delta\big{]}=\frac{\ell}{m+1},\ \ \mathbf{Var}\big{[}\delta\big{]}=\frac{\ell(m+1-\ell)}{(m+1)^{2}(m+2)},$ (2)
that shows by appropriately tuning $m$ and $\ell$, we can significantly reduce the variance $\mathbf{Var}[\delta]$, to avhieve tighter bounds. For instance, when $m=8000$ and $\ell=7999$, the variance of the corresponding Beta distribution is $3.123\times 10^{-8}$, which is extremely small.
However, to present this in a completely formal way, we only considers low variance scenarios ($\mathbf{Var}(\delta)\ll 1$) and we also include this source of uncertainty by reformulating this guarantee with a double-step probablistic guarantee. To include this uncertainty, we can utilize the cumulative density function, CDF, of beta distribution- a function known as regularized incomplete beta function- and propose the following guarantee based on the definition of the CDF.
$\Pr\left[\ \ \Pr[\ \ R^{\text{unseen}}\leq R_{\ell}\ \ ]>1-\epsilon\ \ \right]>1-\mathsf{betacdf}_{1-\epsilon}(\ell,m+1-\ell).$ (3)
where $\mathsf{betacdf}_{1-\epsilon}(\ell,m+1-\ell)$ is the regularized incomplete beta function that is characterized by the choice of $m$ and $\ell$ evaluated at point $\delta_{1}=1-\epsilon$. The one-step marginal guarantee in (1) holds only when the rank is set to $\ell:=\lceil(m+1)(1-\epsilon)\rceil$. However, as shown in Appendix A, a key advantage of the double-step probabilistic guarantee is its flexibility: the strict relationship among $m$, $\ell$, and $\epsilon$ can be relaxed. Therefore, in this paper, we focus on the double-step guarantee and do not enforce the constraint $\ell:=\lceil(m+1)(1-\epsilon)\rceil$. Instead, we tune $m$, $\ell$, and $\epsilon$ such that $\mathsf{betacdf}_{1-\epsilon}(\ell,m+1-\ell)$ becomes negligibly small. Henceforth, for any logical statement $P(\ell,m)\in\left\{\mathrm{true},\mathrm{false}\right\}$, defined over the hyper-parameters $m,\ell\leq m$, we refer to the double-step guarantee,
$\Pr\left[\ \ \Pr[\ \ P(\ell,m)\ \ ]>1-\epsilon\ \ \right]>1-\mathsf{betacdf}_{1-\epsilon}(\ell,m+1-\ell)$
as the $\langle\epsilon,\ell,m\rangle$ guarantee and we say the statement $P(\ell,m)$ satisfies a $\langle\epsilon,\ell,m\rangle$ guarantee. We also refer to $\epsilon$ as the miscoverage level, $\delta_{1}:=1-\epsilon$, as the coverage level and $\delta_{2}:=1-\mathsf{betacdf}_{1-\epsilon}(\ell,m+1-\ell)$ as the confidence of guarantee. Appendix B provides a clarifying example on this topic.
Determinitic and Probabilistic Reachability Analysis on Neural Networks
Given a neural network $f:\mathbb{R}^{n_{0}}\rightarrow\mathbb{R}^{n}$ and an input set $\mathbf{I}\subset\mathbb{R}^{n_{0}}$, deterministic reachability analysis aims to construct a set $\mathbf{R}_{f}(\mathbf{I})\subset\mathbb{R}^{n}$ such that
$x\in\mathbf{I}\quad\Rightarrow\quad f(x)\in\mathbf{R}_{f}(\mathbf{I}).$ (4)
In contrast, probabilistic reachability analysis assumes a distribution $\mathcal{W}$ over the input set $\mathbf{I}$, denoted $x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}$, and for a given miscoverage level $\epsilon$, proposes a set $\mathbf{R}_{f}^{\epsilon}(\mathcal{W})\subset\mathbb{R}^{n}$ such that
$x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}\quad\Rightarrow\quad\Pr\left[f(x)\in\mathbf{R}_{f}^{\epsilon}(\mathcal{W})\right]\geq 1-\epsilon.$ (5)
In this paper, we detail how to compute a probabilistic reach set with miscoverage level $\epsilon$, and sampling distribution $x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}$ by constructing a suitable calibration set $\mathbf{M}$ of size $m$, and selecting an appropriate rank $\ell$. Therefore to keep the terminlogies consistent, we reformulate the probabilistic reach set in (5) using the $\langle\epsilon,\ell,m\rangle$ guarantee, and we denote it by $\mathbf{R}_{f}^{\epsilon}(\mathcal{W}\ ;\ell,m)$, which satisfies
$x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}\quad\Rightarrow\quad\Pr\left[\Pr\left[f(x)\in\mathbf{R}_{f}^{\epsilon}(\mathcal{W}\ ;\ell,m)\right]\geq 1-\epsilon\right]\geq 1-\mathsf{betacdf}_{1-\epsilon}(\ell,m+1-\ell).$
Adversarial Examples and Robustness of SSNs
An adversarial attack on an image involves perturbing the input using a set of noise images $x^{\text{noise}}_{1},\ldots,x^{\text{noise}}_{r}$ and corresponding coefficients $\lambda=[\lambda(1),\ldots,\lambda(r)]^{\top}$. These perturbations are applied through a parameterized function $\Delta_{\lambda,x^{\text{noise}}}(\cdot)$, generating the adversarial image as:
$x^{\text{adv}}=\Delta_{\lambda,x^{\text{noise}}}(x)=x+\sum_{i=1}^{r}\lambda(i)x_{i}^{\text{noise}}.$ (6)
We focus on evaluating the robustness of semantic segmentation networks (SSNs) under such adversarial settings, particularly under the unknown, bounded adversarial examples ($x^{\text{adv}}\in\mathbf{I}$) where the coefficients in vector $\lambda$ are unknown but bounded within specified ranges, i.e., $\lambda\in[\underline{\lambda},\bar{\lambda}]\subset\mathbb{R}^{r}$.
A pixel $x(i,j)$ is deemed attacked if it faces perturbation. It is also considered robust if, for all adversarial examples, the predicted label remains unchanged:  $\mathsf{SSN}(\Delta_{\lambda,x^{\text{noise}}}(x))(i,j)=\mathsf{SSN}(x)(i,j).$
Consider another scenario where there exists perturbations such that the label at $(i,j)$ changes. In this case, the pixel is non-robust if for all perturbations we have $\mathsf{SSN}(\Delta_{\lambda,x^{\text{noise}}}(x))(i,j)\neq\mathsf{SSN}(x)(i,j)$. The pixel is also unknown if for some perturbations we have $\mathsf{SSN}(\Delta_{\lambda,x^{\text{noise}}}(x))(i,j)=\mathsf{SSN}(x)(i,j)$, and for some other perturbations we have $\mathsf{SSN}(\Delta_{\lambda,x^{\text{noise}}}(x))(i,j)\neq\mathsf{SSN}(x)(i,j)$.
To quantify robustness of SSN we consider two different metrics:
Robustness Value.: The Robustness Value (RV) of a network is the percentage of pixels that remain robust under attack: $\quad RV=100\times(N_{\text{robust}}/N_{\text{pixels}}),\quad\text{where }N_{\text{pixels}}=h\times w$.
Problem Formulation
Given a SSN, $x\mapsto\mathsf{SSN}(x)$, we address two problems:
Problem 1. Given an input image $x$, and a desired miscoverage level $\epsilon$, the goal is to determine whether each pixel $x(i,j)$ is robust, non-robust or unknown to a set of adversarial examples $x^{\text{adv}}\in\mathbf{I}$ with strong probabilistic guarantees.
Problem 2. Given $K$ test images $\{x_{1},\ldots,x_{K}\}$, and a set of adversarial examples $\mathbf{I}$, considering a miscoverage level $\epsilon$, compute the average robustness value $\overline{RV}$, with strong probabilistic guarantees.
3 Scaling Probabilistic Reachability Analysis on SSN with Strong Guarantees
Given a specific distribution $x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}$ for sampling inputs $x$ from the input set $\mathbf{I}$, reasoning about the resulting output distribution $y\sim\mathcal{Y}$, is infeasible due to the high degree of nonlinearity in neural networks. As a result, providing probabilistic coverage guarantees over the network’s output space becomes a significant challenge. A key advantage of conformal inference is its distributional robustness—it produces valid guarantees that are robust to all family of distributions that can well represent the samples collected in the calibration dataset. This property makes conformal inference particularly well-suited for reachability analysis of neural networks, which are known for producing complex and often intractable output distributions.
However, a major challenge arises when applied to high-dimensional datasets: The scalar elements $R$ in calibration dataset are defined over the outputs $y\sim\mathcal{Y}$ and the space of all possible distributions $\mathcal{Y}$ that can adequately represent the calibration dataset is significantly larger in higher dimensions and thus conformal inference tends to be overly conservative. This poses a significant obstacle when applying CI to semantic segmentation networks (SSNs), which are known for producing extremely high-dimensional outputs. In this paper, we propose a novel learning-based neural network reachability analysis method for SSNs, that leverages deflative principal component analysis (PCA) (Mackey, 2008) to address the dimensionality challenges inherent in conformal inference when applied to provide coverage guarantees on neural network with high dimensional outputs.
In this section, we present our reachability technique in an incremental manner to clearly illustrate both the contributions and the challenges involved in applying conformal inference to SSNs. We begin by introducing a naive baseline approach, then highlight its limitations on high-dimensional spaces. Subsequently, we describe the enhancements we propose to overcome these challenges, culminating in the final version of our reachability algorithm.
3.1 Naive Reachability Technique via Conformal Inference
The first step to find a probabilistic reachset on neural networks with conformal inference (CI) is to generate a calibration dataset $\mathbf{M}$ of size $m$ and a training dataset $\mathbf{T}$ of size $t$.
Generating train dataset. We sample $t$ different inputs $x_{j}^{\text{train}},j=1,2,\ldots,t$ from $\mathbf{I}$, with any distribution of interest, $x\stackrel{{\scriptstyle\mathcal{W^{\prime}}}}{{\sim}}\mathbf{I}$, and compute their corresponding outputs $\mathbf{vec}(y_{j}^{\text{train}})=f(\mathbf{vec}(x_{j}^{\text{train}}))$. This gives us the training dataset, $\mathbf{T}=\left\{(x_{1}^{\text{train}},y_{1}^{\text{train}}),(x_{2}^{\text{train}},y_{2}^{\text{train}}),\ldots,(x_{t}^{\text{train}},y_{t}^{\text{train}})\right\}$.
Generating Calibration dataset. We sample the input dataset with $m$ different inputs $x^{\text{calib}}_{i},i=1,2,\ldots,m$ from the distribution $x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}$ and their corresponding set of outputs $\mathbf{vec}(y^{\text{calib}}_{i})=f(\mathbf{vec}(x^{\text{calib}}_{i})),i=1,2,\ldots,m$. This gives us the input/output dataset $\mathbf{IO}=\left\{(x_{1}^{\text{calib}},y^{\text{calib}}_{1}),(x_{2}^{\text{calib}},y^{\text{calib}}_{2}),\ldots,(x_{m}^{\text{calib}},y^{\text{calib}}_{m})\right\}$. We next attempt to use $\mathbf{IO}$ to compute a suitable nonconformity score, $R\in\mathbb{R}_{\geq 0}$. To that end, inspired by recent work on using conformal inference in time-series Cleaveland et al. (2024), we design the nonconformity scores as follows in Equation (7). This design provides us a hyper-rectanglular set. Assuming the output $\mathbf{vec}(y)=\left[\mathbf{vec}(y)(1),\ldots,\mathbf{vec}(y)(n)\right]\in\mathbb{R}^{n}$, we propose, the nonconformity scores $R^{\text{calib}}_{i}$ as,
$R^{\text{calib}}_{i}=\max\left(\frac{|\mathbf{vec}(y^{\text{calib}}_{i})(1)-c(1)|}{\tau_{1}},\ldots,\frac{|\mathbf{vec}(y^{\text{calib}}_{i})(n)-c(n)|}{\tau_{n}}\right),\quad i=1,2,\ldots,m$ (7)
where $c\in\mathbb{R}^{n}$ is the average of the members in the train dataset $\mathbf{T}$, and for $k=1,2,\ldots,n$, the normalization factor $\tau_{k}$ is proposed to normalize the elements of the random vectors $|\mathbf{vec}(y)-c|$ which can also be obtained using the train dataset as,
$\tau_{k}:=\max\left(\ \ \tau^{*},\max\left(\ \ \ |\mathbf{vec}(y_{1}^{\text{train}})(k)-c(k)|,\ \ \ldots,\ \ |\mathbf{vec}(y_{t}^{\text{train}})(k)-c(k)|\ \ \ \right)\ \ \right),$ (8)
where we set $\tau^{*}=10^{-5}\left(\sum_{k=1}^{n}\sum_{j=1}^{t}|\mathbf{vec}(y_{j}^{\text{train}})(k)-c(k)|\right)/nt$, to avoid facing zero as our normalization factor. We then propose the calibration dataset as $\mathbf{M}=\left\{R^{\text{calib}}_{1},R^{\text{calib}}_{2},\ldots,R^{\text{calib}}_{m}\right\}$.
The next step is to sort the nonconformity scores in $\mathbf{M}$ based on their magnitude. Without loss of generality let’s assume $R^{\text{calib}}_{1}<R^{\text{calib}}_{2}<\ldots<R^{\text{calib}}_{m}$. Given a new draw $x^{\text{unseen}}$ from the distribution $x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}$ and the architecture of the neural network, the output $\mathbf{vec}(y^{\text{unseen}})=f(\mathbf{vec}(x^{\text{unseen}}))$ follows a specific distribution. We call this distribution as $y\sim\mathcal{Y}$. Furthermore given the map we proposed to compute the nonconformity scores, the i.i.d random variables,
$R^{\text{unseen}}=\max\left(\frac{|\mathbf{vec}(y^{\text{unseen}})(1)-c(1)|}{\tau_{1}},\ldots,\frac{|\mathbf{vec}(y^{\text{unseen}})(n)-c(n)|}{\tau_{n}}\right)$ (9)
adopt another distribution where we call it $R\sim\mathcal{D}$. Here $R^{\text{unseen}}$ and the elements $R^{\text{calib}}_{1},R^{\text{calib}}_{2},\ldots,R^{\text{calib}}_{m}$ are all sampled from the distribution $\mathcal{D}$. This means that we can apply conformal inference. Therefore, for the new draw $R^{\text{unseen}}$, given a desired rank, $\ell\leq m$ and a desired miscoverage level $\epsilon$, we can propose the following $\langle\epsilon,\ell,m\rangle$ guarantee,
$\Pr\left[\ \ \Pr[\ \ R^{\text{unseen}}\leq R^{\text{calib}}_{\ell}\ \ ]>1-\epsilon\ \ \right]>1-\mathsf{betacdf}_{1-\epsilon}(\ell,m+1-\ell).$
Following the definition in Equation (9), the logical statement $P_{1}(\ell,m):=[R^{\text{unseen}}\leq R_{\ell}^{\text{calib}}]$ is logically equivalent to
$P_{2}(\ell,m):=\bigwedge_{k=1}^{n}\left[c(k)-\sigma_{k}\leq\mathbf{vec}(y^{\text{unseen}})(k)\leq c(k)+\sigma_{k}\right],\quad\text{where}\ \sigma_{k}=\tau_{k}R^{\text{calib}}_{\ell},$ (10)
which defines a hyper-rectangular set as the reachset for any unseen output of the neural network generated by $x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\ \mathbf{I}$. Since $P_{1}(\ell,m)$ satisfies the $\langle\epsilon,\ell,m\rangle$ guarantee, this reachset also provides the same coverage guarantee on the unseen outputs of the neural network. Appendix C provides a clarifying example on this technique demonstrating that it is effective even on very deep neural networks, offering strong probabilistic guarantees with a reasonable number of required sampling. As the architecture of a deep neural network becomes more complex, the distribution of its output can exhibit increasingly intricate behavior. Conformal inference remains robust to any output distribution that represents well the calibration dataset, allowing it to handle such complexities reliably. However, the tightness of the reachset does not persist when the dimensionality of the network’s output increases.
  • The first issue is that the naive technique can only produce a hyper-rectangular reachset, that can be a significant source of conservative in high-dimensional spaces.
  • The second issue is that the space of distributions $\mathcal{Y}$ that can provide some distribution $\mathcal{D}$ which well represent the calibration dataset is significantly larger in high dimensional space.
3.2 From Hyper-Rectangular Reachsets to Convex Hull Representations
Our approach to constructing a general convex hull as the reachset is based on approximating the original deep neural network $f$ with another computation graph $g:\mathbb{R}^{n_{0}}\to\mathbb{R}^{n}$, for which a guaranteed deterministic reachable set can be derived. The model $g$ is generated based on the training dataset $\mathbf{T}$, which implies it is accurate only for inputs $x$ sampled from the set $\mathbf{I}$. The core idea is to utilize its deterministic reachable set, and then inflate it to account for approximation error between $f$ and its surrogate $g$, thereby obtaining a probabilistic reachset for the original model $f$ with an $\langle\epsilon,\ell,m\rangle$ guarantee.
Obtaining the surrogate model $g$ serves a dual purpose: it not only provides an approximation for $f$, but also defines the shape of a convex hull which when inflated serves as a tighter over-approximation of the probabilistc reachset of $f$. Once $g$ is computed, this convex shape can be computed as a convex hull. To inflate this surrogate reachset, we follow the procedure outlined in the naive technique (Section 3.1) to generate a hyper-rectangle that bounds the prediction error with a provable $\langle\epsilon,\ell,m\rangle$ guarantee. In other words, we define $q(\mathbf{vec}(x))=f(\mathbf{vec}(x))-g(\mathbf{vec}(x))$ and apply the naive technique on $q(\mathbf{vec}(x))$ to compute its hyper-rectangular reachset with $\langle\epsilon,\ell,m\rangle$ guarantee. We refer to this hyper-rectangle as inflating set.
In conclusion, if we denote the deterministic reachset of surrogate model as the starset $\mathcal{R}_{g}(\mathbf{I})$,
$x\in\mathbf{I}\ \ \Rightarrow\ \ g(\mathbf{vec}(x))\in\mathcal{R}_{g}(\mathbf{I})$
and for some hyper-parameters $(\ell,m)$ and distribution $x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}$, the $\langle\epsilon,\ell,m\rangle$ guaranteed probabilistic reachset for prediction errors by $\mathcal{R}_{q}^{\epsilon}(\mathcal{W};\ \ell,m)$,
$x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}\ \Rightarrow\ \Pr\left[\ \Pr[\ \ q(\mathbf{vec}(x))\in\mathcal{R}_{q}^{\epsilon}(\mathcal{W};\ \ell,m)\ ]>1-\epsilon\ \right]>1-\mathsf{betacdf}_{1-\epsilon}(\ell,m+1-\ell),$
then the probabilistic reachset of the model $f$ with the same $\langle\epsilon,\ell,m\rangle$ guarantee is obtainable by $\mathcal{R}_{f}^{\epsilon}(\mathcal{W};\ell,m)=\mathcal{R}_{g}(\mathbf{I})\oplus\mathcal{R}_{q}^{\epsilon}(\mathcal{W};\ \ell,m)$, where $\oplus$ denotes the Minkowski sum,
$x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}\ \ \Rightarrow\ \ \Pr\left[\ \ \Pr[\ \ f(\mathbf{vec}(x))\in\mathcal{R}_{f}^{\epsilon}(\mathcal{W};\ell,m)\ \ ]>1-\epsilon\ \ \right]>1-\mathsf{betacdf}_{1-\epsilon}(\ell,m+1-\ell).$
This technique offers two key improvements over the naive approach presented in Section 3.1:
  • The reachset is no longer constrained to a hyper-rectangle, eliminating the first source of conservatism inherent in the naive method.
  • The calibration dataset is now defined using prediction errors rather than the network’s raw outputs. Since prediction errors are typically of much smaller magnitude than the outputs of model $f$, this significantly reduces the conservatism of conformal inference in high-dimensional spaces. This implies that a useful surrogate model must maintain an acceptable level of fidelity.
To obtain a surrogate model $g$ with high fidelity while ensuring a deterministic reachable set, we propose using the convex hull $\mathbf{CH}$ over the training outputs $y_{j}^{\text{train}},\;j=1,\dots,t$,
$\mathbf{CH}=\sum_{j=1}^{t}\alpha_{j}\mathbf{vec}(y_{j}^{\text{train}}),\qquad\mathbf{s.t.}\quad\sum_{j=1}^{t}\alpha_{j}=1,\quad\alpha_{1},\cdots,\alpha_{t}>0$
and then incorporating a clipping block (see Appendix D) that projects the outputs of $f$ for any unseen image $x^{\text{unseen}}$, into this convex hull. Denoting the clipping block with $\mathbf{CLP}$ we propose the surrogate model as: $g(\mathbf{vec}(x))=\mathbf{CLP}(f(\mathbf{vec}(x)))$, that makes certain $\mathbf{R}_{g}(\mathbf{I})=\mathbf{CH}$.
Despite the improvements introduced in this section, reachability analysis remains unscalable in high-dimensional settings. The fundamental difficulty lies in the fact that projecting $\mathbf{vec}(y^{\text{unseen}})$ on $\mathbf{CH}$ does not scale well with the dimensionality of the output space. To mitigate this, we incorporate an scalable version of Principal Component Analysis (PCA), known as defelation algorithm, before clipping, a technique known to be effective for handling PCA in high-dimensional spaces. Therefore, the error $q(x)$ primarily arises from the inclusion of PCA and clipping block.
Initialize $z_{1}\ ,\ z_{2},\ \ldots,\ z_{t}\ \leftarrow\ \mathbf{vec}(y_{1}^{\text{train}}),\ \mathbf{vec}(y_{2}^{\text{train}}),\ \ldots,\ \mathbf{vec}(y_{t}^{\text{train}})$$A\leftarrow\ \left[\ \right]$
for $\mathsf{index}=0,1,\ldots N-1$ do
 // $\mathtt{Train\ the\ principal\ direction\ using\ stochastic\ gradient\ ascent}$
 $\vec{a}_{\mathsf{index}}\ \ \leftarrow\ \ \displaystyle\arg\max_{\vec{a}}$$\left[\mathcal{J}:=\frac{1}{t}\sum_{j=1}^{t}\vec{a}^{\top}z_{j}z_{j}^{\top}\vec{a}\right],\qquad\mathbf{s.t.}\quad\|\vec{a}\|_{2}=1$
 $A.\mathbf{append}(\vec{a}_{\mathsf{index}})$
    // $\mathtt{collect\ the\ principal\ direction\ in\ A}$
    // $\mathtt{Update\ the\ dataset\ by\ component\ removal\ along\ principal\ direction.}$
 $z_{1},\ z_{2},\ \ldots,\ z_{t}\ \leftarrow\ z_{1}-(\vec{a}_{\mathsf{index}}^{\top}z_{1})\vec{a}_{\mathsf{index}},\ z_{2}-(\vec{a}_{\mathsf{index}}^{\top}z_{2})\vec{a}_{\mathsf{index}},\ \ldots,\ z_{t}-(\vec{a}_{\mathsf{index}}^{\top}z_{t})\vec{a}_{\mathsf{index}}$
end for
Algorithm 1 Learning Based Principal Component Analysis through Deflation Algorithm
3.3 Surrogate Model for SSNs via Principal Component Analyis
In this section, we present our methodologies to mitigate the dimensionality challenges to provide the surrogate model, and present our technique for the robustness analysis of SSNs.
In our robustness analysis of SSNs, as indicated by equation (6), adversarial perturbations are confined to an $r$-dimensional subspace of the input space. Thus, we reformulate the surrogate as,
$\vskip-5.69054ptg^{\prime}(\lambda)=g(\ \mathbf{vec}(\ x+\sum_{i=1}^{r}\lambda(i)x^{\text{noise}}\ )\ )=g(\mathbf{vec}(x^{\text{adv}}))\quad\text{and}\quad\ \lambda\in[\underline{\lambda},\overline{\lambda}]\subset\mathbb{R}^{r},$
and also the deterministic reachset $\mathbf{R}_{g}(\mathbf{I})$ based on the vector $\lambda$ as $\mathbf{R}_{g^{\prime}}([\underline{\lambda},\overline{\lambda}])=\mathbf{R}_{g}(\mathbf{I})=\mathbf{CH}$.
High-Dimensionality of Output Space. To address this challenge, we firt reduce the dimensionality of the logits i.e., for a choice of $N\ll n$, and then generate the convex hull to be used in the clipping process. Let’s sample training images $x_{j}^{\text{train}}$, for $j=1,\ldots,t$, from the adversarial set $\mathbf{I}$, using sampling the coefficient vectors $\lambda_{j}^{\text{train}}$ form $\underline{\lambda},\overline{\lambda}]$. Then the corresponding logits $\mathbf{vec}(y_{j}^{\text{train}})$ form a point cloud in $\mathbb{R}^{n}$. This stage aims to train the top $N$ principal directions of cloud, which are stored as columns of the matrix $A=[\vec{a}_{0},\ \vec{a}_{1},\ \ldots,\ \vec{a}_{N-1}]\in\mathbb{R}^{n\times N}$.
We utilize this matrix to represent the high dimensional logits $y_{j}^{\text{train}}$ with a lower dimentional representative $v_{j}\in\mathbb{R}^{N}$ where $v_{j}=A^{\top}\mathbf{vec}(y_{j}^{\text{train}})$. Principal Component Analysis is a technique to reduce the dimensionalty, but it does not scale to high dimensional spaces. To address this issue, deflation algorithms have been proposed (see Algorithm 1) that compute principal components iteratively, in order of significance (Mackey, 2008). At each iteration, once the most dominant principal component is identified, the dataset is projected onto the orthogonal complement of that direction—effectively removing its contribution—before computing the next principal direction.
In this paper, we adopt a learning based deflation algorithm that is presented in Algorithm 1. In this algorithm we set the components of the principal direction as trainable parameters and maximize the variance of the dataset in the direction of this vector. At each iteration of this algorithm, the optimization to train the principal direction have one global maxima one global minima and n2n-2 saddle points, which $n$ is the dimension of logits $\mathbf{vec}(y_{j}^{\text{train}})$ Mackey (2008). Therefore, although the training operates in a high-dimensional space, convergence to global maxima is guaranteed.
Next we generate the convex hull over the cloud of point of the low dimensional representatives of logit $v_{j},j=1,\ldots,t$ and then apply the clipping block. Thus, for any unseen $\lambda$ we define,
$g^{\prime}(\lambda^{\text{unseen}})=g(x^{\text{adv}})=A\ \mathbf{CLP}(A^{\top}f(\mathbf{vec}(\ x+\sum_{i=1}^{r}\lambda^{\text{unseen}}(i)x^{\text{noise}}\ )))$
In conclusion we present the surrogate model for $f$ as $g(x^{\text{adv}})$ which is a scalable choice for SSNs.
Detecting Robust, Nonrobust and Unknown Pixels with Strong Probabilistic Guarantees
Once the reachset over the SSN logits $y\in\mathbb{R}^{n}$ is constructed as a starset, we project it onto each logit component. This allows us to determine the range of values for each class label $l\in\mathbf{L}$ at every pixel location $(i,j)\in\{1,\ldots,h\}\times\{1,\ldots,w\}$. Consequently, each pixel $(i,j)$ is associated with $L$ logit intervals, one for each class. Let $l^{*}\in\mathbf{L}$ denote the class whose logit interval has the highest lower bound at pixel $(i,j)$. If this lower bound is not strictly greater than the upper bounds of all other classes’ intervals, the pixel is labeled as unknown as it can obtain more than one labels under UBAA. In case it is greater, then if $l^{*}=\mathsf{SSN}(x)(i,j)$, the prediction at that pixel is labeled as robust and if $l^{*}\neq\mathsf{SSN}(x)(i,j)$, it is labeled as non-robust. The Algorithm 2 explains this procedure in detail.
4 Experiments
In our numerical evaluation, we pose three different research questions and address each of them using numerical results. The main one is available in this section and the rest of them are in Appendix E. We utilized a Linux machine, with $48$ GB of GPU memory, $512$ GB of RAM, and $112$ CPUs.
Input: $\mathbf{I}$, $x$, $f$, $\mathcal{W}$, $\epsilon$, Hyper-parameters(,m)($\ell$,m)
Output: The status of pixels
// $\mathtt{Project\ the\ Starset}$
$\underline{y},\overline{y}]\leftarrow\mathcal{R}_{f}^{\epsilon}(\mathcal{W};\ell,m)$
foreach $(i,j)\in\{1{:}h\}\times\{1{:}w\}$ do
 $l^{*}\leftarrow\arg\max_{l}\underline{y}(i,j,l)$
 if $\underline{y}(i,j,l^{*})\leq\max_{l\neq l^{*}}\overline{y}(i,j,l)$ then
      Label pixel $(i,j)$ as unknown
   end if
 else if $l^{*}=\mathsf{SSN}(x)(i,j)$ then
      Label pixel $(i,j)$ as robust
   end if
 else
      Label pixel $(i,j)$ as nonrobust
   end if
end foreach
return Pixel-wise labeling
Algorithm 2 Detection of the pixel status
Figure 1: Illustrates the perturbation set $\mathbf{I}$ corresponding to an $r$-dimensional darkening adversary on image $x$. This set is made by independent perturbations on all $nc$ channels in $r^{\prime}$ different pixels ($r=nc\times r^{\prime}$), each with R,G, and B intensities greater than $150/255$. The lower bound $\mathbf{LB}$ represents the maximum darkening (channel intensities set to zero), while the upper bound $\mathbf{UB}$ corresponds to minimum darkening.
Figure 2: Shows $\overline{\mathbf{RV}}$ versus perturbation level and perturbation dimension. In the top row, we perturbed the entire image with $\ell_{2}$ perturbation for CamVid and Lung Segmentation and $\ell_{\infty}$ for OCTA-500. In the bottom row, we consider darkening adversary with $e=5/255$ and we increase the dimension of perturbation. Robustness is averaged over 200 test images .
Table 1: Model details and probabilistic guarantees. Given hyperparameters $\ell$, $m$, and coverage $\delta_{1}=1-\epsilon$, the confidence $\delta_{2}$ is computed for each experiment. The reported verification runtime is averaged over all experiments run on a model. One important point is, our verification runtime depends mainly on inference time, and hyperparameters ($\ell,m$), and is only slightly sensitive to perturbation magnitude and dimension $(e,r)$. For per-experiment runtimes, see Figure 5 in Appendix E.
Dataset Model Input Output Number of $(m,m-\ell)$ coverage confidence Average
name name dimension dimension Parameters $\delta_{1}$ $\delta_{2}$ runtime (method)
Lung Segmentation UNet1 $\times\!512\!\times\!1 512×512$$\times\!512\!\times\!1$ $14,779,841$ $(8\mathbf{e}3,1)$ 0.999 0.997 5.6 min (Surrogate)
OCTA-500 UNet2 $\times\!304\!\times\!1 304×304$$\times\!304\!\times\!1$ $5,478,785$ $(8\mathbf{e}3,1)$ 0.999 0.997 9.8 min (Surrogate)
CamVid BiSeNet $720\!\times\!960\!\times\!3$ $720\!\times\!960\!\times\!12$ $12,511,084$ $(8\mathbf{e}3,1)$ 0.999 0.997 18.3 min (Surrogate)
Cityscapes (Appendix E) HRNetV2 $1024\!\times\!2048\!\times\!3$ $256\!\times\!512\!\times\!19$ $65,859,379$ $(921,1)$ 0.99 0.997 3.5 min (Naive)
RQ1: How does the technique scale with model size, number of perturbed pixels, and perturbation level? How does output high dimensionality, and the irregularity in output distribution of large and complex NNs make our guarantees overly conservative?
To assess scalability, we test our verification method on high-dimensional datasets (e.g., Lung Segmentation (Jaeger et al., 2014; Candemir et al., 2014), OCTA-500 (Li et al., 2019), CamVid (Brostow et al., 2009)) using large pre-trained models. We evaluate performance on a subset of $200$ test images across varying perturbation dimensions and magnitudes, with results shown in Figure 2. Model details and probabilistic guarantees are summarized in Table 1, where we show the following $\langle\epsilon,\ell,m\rangle$ guarantee:
[…]
In these experiments, we studied an $r$-dimensional darkening adversary (see Figure 1), where $r^{\prime}$ pixels in an image $x$ with intensity above $150/255$ in all channels are randomly selected for perturbation ($r=nc\times r^{\prime}$). Each direction $x_{i}^{\text{noise}}$ corresponds to darkening a channel of one such pixels, and $x^{\text{adv}}$ is parameterized by an $r$-dimensional coefficient vector $\lambda\in[\underline{\lambda},\overline{\lambda}]$. Here, $\overline{\lambda}$ induces full darkening (intensity zero), while $\underline{\lambda}$ applies partial darkening. This bound defines the perturbation space $x^{\text{adv}}\in\mathbf{I}$. We also present a demo for the status of all pixels in the segmentation mask in Figures 6,7 and 8.
To assess the conservatism, we examine the projection bounds $\underline{y},\overline{y}]$ introduced in Algorithm 2. Specifically, we sample $10^{6}$ adversarial examples from $x^{\text{adv}}\stackrel{{\scriptstyle\scriptscriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}$ and use them to report:
  • Empirical miscoverage $\hat{\epsilon}$, that is the percentage of events, where $f(\mathbf{vec}(x^{\text{adv}}))\notin[\underline{y},\overline{y}]$.
  • Empirical bound $\underline{\hat{y}},\overline{\hat{y}}$], via component-wise minima/maxima of $f(\mathbf{vec}(x^{\text{adv}}))$ across samples.
The degree of conservatism can be assessed by comparing $\underline{\hat{y}},\overline{\hat{y}}$] with $\underline{y},\overline{y}]$. To measure this, we compute $\mathsf{bound\_ratio}=\sum_{k=1}^{h\times w\times L}(\overline{\hat{y}}(k)-\underline{\hat{y}}(k))/\sum_{k=1}^{h\times w\times L}(\overline{y}(k)-\underline{y}(k))$ and due to the high cost of $10^{6}$ inferences on the models, we only perform the conservatism analysis on one specific case with BiSeNet from Figure 2, Table 2 in Appendix E details the numerical results.
5 Conclusion
Our probabilistic verification method is both scalable and data-efficient. While it does not offer deterministic guarantees, it can provide meaningful assurances in regions where existing deterministic tools are not applicable due to computational challenges. In this paper, we also demonstrated this capability through a series of numerical experiments.
Appendix A Double-Step Probabilistic Guarantee
The necessary steps to establish the two-sided probabilistic guarantee, along with the flexibility in choosing the dataset size $m$, rank $\ell$, and failure probability $\epsilon$, begin with the theory of order statistics David and Nagaraja (2004). This flexibility means that the double-step probability constraint remains valid regardless of the particular values chosen for $m$, $\ell\leq m$, and $\epsilon\in[0,1]$. The following lemma introduces the concept of order statistics.
Lemma 1 (From David and Nagaraja (2004))
Consider $m$ independent and identically distributed (i.i.d.), real-valued data points $x\in[0,1]$ drawn from uniform distribution $x\sim\mathcal{U}$. Suppose we sort them in ascending order and denote the $\ell^{th}$ smallest number by $x_{\ell}$,(i.e., we have $x_{1}<x_{2}<\ldots<x_{m}$). Let $\mathbf{Beta}(\alpha,\beta)$ denote the Beta distribution [^3] The Beta distribution is a family of continuous probability distributions defined on the interval $0\leq x\leq 1$ with shape parameters $\alpha$ and $\beta$, and with probability density function $f(x;\alpha,\beta)=\frac{x^{\alpha-1}(1-x)^{\beta-1}}{B(\alpha,\beta)}$, where the constant $B(\alpha,\beta)=\frac{\Gamma(\alpha)\Gamma(\beta)}{\Gamma(\alpha+\beta)}$ and $\Gamma(z)=\int_{0}^{\infty}t^{z-1}e^{-t}\mathbf{d}t$ is the Gamma function.. Then the uniform random variable at rank $\ell$ follows:
$x_{\ell}\sim\mathbf{Beta}(\ell,m+1-\ell)$
In the context of conformal inference Lemma 1, has been extended from a uniform and domain-bounded distribution to a general distribution as follows:
Lemma 2 (From Vovk et al. (2005))
Consider $m$ independent and identically distributed (i.i.d.), real-valued data points $R$ drawn from some distribution $R\sim\mathcal{D}$. Suppose we sort them in ascending order and denote the $\ell^{th}$ smallest number by $R_{\ell}$, (i.e., we have $R_{1}<R_{2}<\ldots<R_{m}$). Let $\mathbf{Beta}(\alpha,\beta)$ denote the Beta distribution. Then, for an arbitrary $R_{m+1}$ drawn from the same distribution $\mathcal{D}$, the following holds:
$\Pr\left[R_{m+1}<R_{\ell}\right]\sim\mathbf{Beta}(\ell,m+1-\ell),\quad 1\leq\ell\leq m.$ (11)
This result is motivated by two key observations: (1) for a continuous distribution $R\sim\mathcal{D}$, the cumulative probability $\Pr[R\leq r]$ is uniformly distributed on $[0,1]$, and (2) sorting the samples $R_{1}<R_{2}<\ldots<R_{m}$ implies that the corresponding cumulative probabilities $\Pr[R\leq R_{\ell}]$ are also ordered, i.e., $\Pr[R\leq R_{1}]<\Pr[R\leq R_{2}]<\ldots<\Pr[R\leq R_{m}]$. These observations implies that for all $\ell=0,1,2,\ldots,m$, the random variable $\delta:=\Pr[R\leq R_{\ell}]$ satisfies the conditions outlined in Lemma 1.
In simple terms, this means that while a fixed dataset $R_{1},R_{2},\ldots,R_{m}$ is not sufficient to determine the exact value of $\delta=\Pr[R<R_{\ell}]$, we can infer that the value of $\delta$ must follow a Beta distribution that is solely defined based on $m$ and $\ell$. This allows us to tune the parameters $m$ and $\ell$ to provide strong guarantees for the coverage $R<R_{\ell}$ where $R_{\ell}$ is obtained from our fixed sampled dataset.
This guarantee can be equivalently expressed as a two-sided probabilistic bound. Specifically, let $\mathsf{betacdf}_{\delta}(\ell,m+1-\ell)$ denote the cumulative distribution function (CDF) of the Beta distribution—also known as the regularized incomplete beta function—evaluated at $\delta$. Then, based on the definition of cumulative density function, CDF, for any confidence $\delta_{1}\in[0,1]$ and any arbitrary distribution $R\sim\mathcal{D}$, the following inequality holds:
$\Pr\left[\ \ \Pr[\ \ R\leq R_{\ell}\ \ ]>\delta_{1}\ \ \right]>1-\mathsf{betacdf}_{\delta_{1}}(\ell,m+1-\ell).$
Note that the key assumption is that all variables $R,R_{1},R_{2},\ldots,R_{m}$ are independently drawn from the same distribution $\mathcal{D}$, and that the previously sampled values are sorted such that $R_{1}<R_{2}<\ldots<R_{m}$. Importantly, the resulting probabilistic guarantee remains valid for any arbitrary choice of sample size $m$, rank $\ell$, and confidence level $\delta_{1}$, provided they lie within appropriate bounds.
Appendix B A Clarifying Example for Conformal Inference
Assume an arbitrary distribution $R\sim\mathcal{D}$. We sample a calibration dataset of size $m=100,000$, we sort the samples as $R_{1}<R_{2}<\ldots<R_{100,000}$ and we select the sample at rank $\ell=99,999$. In this case, given an unseen draw $R^{\text{unseen}}$ from $\mathcal{D}$ for a miscoverage level $\epsilon=0.0001$, the $\langle\epsilon,\ell,m\rangle$ guarantee for logical statement $P(\ell,m)=[R^{\text{unseen}}<R_{\ell}]$ is:
$\Pr\left[\ \ \Pr[\ \ R^{\text{unseen}}<R_{99,999}\ \ ]\ >\ 0.9999\ \ \right]>\ 0.9995008$
which is valid for any calibration dataset $\mathbf{M}=\left\{R_{1},R_{2},\ldots,R_{100,000}\right\}$ sampled from distribution $\mathcal{D}$ if and only if $R^{\text{unseen}}$ is also sampled from $\mathcal{D}$, and $R_{99,999}$ is the $99,999^{\text{th}}$ smallest member of $\mathbf{M}$.
Appendix C A Clarifying Example on Naive Approach
Figure 3: The figure depicts two reachable sets corresponding to different $\langle\epsilon,\ell,m\rangle$ guarantees. Both $S_{1}$ and $S_{2}$ are obtained using the naïve approach, with $S_{2}$ providing a stronger guarantee than $S_{1}$.
Figure 4: The figure compares two reachable sets with the same $\langle\epsilon,\ell,m\rangle$ guarantees. $S_{1}$ is obtained using the naïve approach, while $S_{3}$ is produced by the surrogate-based technique, illustrating the accuracy of the latter.
Since the naive technique is architecture-agnostic we apply that for the reachability analysis of a feedforward $\mathrm{ReLU}$ neural network with $60$ hidden-layers of width $100$, an input layer of dimension $784$ and an output layer of dimension $2$. The set of inputs is $\mathbf{I}:=[0,1]^{784}$ and we generate both calibration and train datasets by sampling $\mathbf{I}$ uniformly. e.g., $\mathcal{W},\mathcal{W}^{\prime}$ are both uniform distributions.
1:  We firstly generate a calibration dataset of size $m=200,000$ and a train dataset of size $t=10,000$, we also consider the rank $\ell=199,998$ and we target the miscoverage level of $\epsilon=0.0001$. In this case we compute a reachable set, $\mathbf{S}_{1}$ with the following $\langle\epsilon,\ell,m\rangle$ guarantee.
$x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}\ \ \Rightarrow\ \ \Pr\left[\ \ \Pr[\ \ f(x)\in\mathbf{S}_{1}\ \ ]>0.9999\ \ \right]>0.9999995.$ (12)
2:  Secondly, we set $m=8464287$, $\ell=8464286$ and $\epsilon=0.000002$ which results in another reachset $\mathbf{S}_{2}$ with the following $\langle\epsilon,\ell,m\rangle$ guarantee.
$x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}\ \ \Rightarrow\ \ \Pr\left[\ \ \Pr[\ \ f(x)\in\mathbf{S}_{2}\ \ ]>0.999998\ \ \right]>0.9999992.$ (13)
Figure 4 shows the reachable sets $\mathbf{S}_{1}$ and $\mathbf{S}_{2}$ in the presence of $10^{6}$ new simulations, $f(x),x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}$, for validation. Our calculations show that the proportion of points lying outside of $\mathbf{S}_{1}$ and $\mathbf{S}_{2}$ are 0.000013 and 0.000001, respectively, which aligns well with the proposed $\langle\epsilon,\ell,m\rangle$ guarantees.
Remark 1
One might argue that the dimensionality issue can be addressed by partitioning the output layer and performing reachability analysis on each partition separately. However, this approach is suboptimal, as the overall probabilistic guarantee deteriorates. Specifically, the combined guarantee is governed by a union bound over the individual partitions, which can lead to an unacceptably loose $\langle\epsilon,\ell,m\rangle$ guarantee.
Appendix D More Details on Clipping Block, Surrogate Model & Minkowski Sum
The clipping block is the core component of our surrogate model, providing both an accurate approximation of the original system and a deterministic reachable set. Specifically, we first generate a cloud of outputs from the training dataset and construct a convex hull enclosing these points. We formulate this convex hull as,
$\mathbf{CH}=\sum_{j=1}^{t}\alpha_{j}\mathbf{vec}(y_{j}^{\text{train}}),\qquad\mathbf{s.t.}\quad\sum_{j=1}^{t}\alpha_{j}=1,\quad\alpha_{1},\cdots,\alpha_{t}>0$
The clipping block then applies a convex optimization procedure that projects every unseen output $\mathbf{vec}(y^{\text{unseen}})$ onto this hull, as $\mathbf{vec}(\hat{y}^{\text{unseen}})=\sum_{j=1}^{t}\alpha^{*}_{j}\mathbf{vec}(y_{j}^{\text{train}})$, where,
$\alpha^{*}_{1},\cdots,\alpha^{*}_{t}=\underset{\alpha_{1},\cdots,\alpha_{t}}{\mathbf{argmin}}\|\mathbf{vec}(y^{\text{unseen}})-\sum_{j=1}^{t}\alpha_{j}\mathbf{vec}(y_{j}^{\text{train}})\|_{l}\quad\mathbf{s.t.}\ \sum_{j=1}^{t}\alpha_{j}=1,\ \ \alpha_{1},\cdots,\alpha_{t}>0,$
that for $l=2$ is a quadratic programming and for $l=1,\infty$ is a linear programming. Among the three options, we are most interested in $l=\infty$, since it minimizes the maximum error across the logit components, making it a suitable choice for reachability analysis. In addition, the $l=2$ formulation is far more memory-intensive compared to the $l=\infty$ case. When we ran it on the OCTA-500 dataset using the Vortex machine (540 GB RAM), it resulted in an out-of-memory error. We also attempted the PGD variant of quadratic programming for $l=2$; however, this led to severe CPU thread oversubscription issues. Thus, the LP formulation scales more gracefully. Each problem is linear, and batching helps amortize model builds, and parallelism is also effective — which is why we observe it running stably on our setup.
Thus, our surrogate model $g$ is composed of two main components. The first is the base model $f$, which takes the vectorized image $\mathbf{vec}(x^{\text{unseen}})$ and produces the corresponding logit vector $\mathbf{vec}(y^{\text{unseen}})$. The second is the clipping block, which takes $\mathbf{vec}(y^{\text{unseen}})$ as input and returns its projection onto the convex hull $\mathbf{CH}$, denoted as $\mathbf{vec}(\hat{y}^{\text{unseen}})$. In other words,
$g(\mathbf{vec}(x^{\text{unseen}}))=\mathbf{vec}(\hat{y}^{\text{unseen}})$
and obviously $\mathbf{CH}$ represents the deterministic reachable set of the surrogate model, $\mathbf{R}_{g}(\mathbf{I})=\mathbf{CH}.$
To generate the convex hull, it is not necessary to include the entire training dataset. Instead, one can identify the extreme points Sartipizadeh and Vincent (2016) and construct the convex hull using only those points, which reduces the number of coefficients $\alpha$ without sacrificing accuracy. This approach makes the convex hull more efficient to be used as it reduces the number of coefficients $\alpha$. Existing algorithms for identifying extreme points include Quickhull Barber et al. (1996) and Clarkson’s algorithm Clarkson (1988). Their computational complexities are $\mathcal{O}(t\log(t))$ for $n=2,3$ and $\mathcal{O}(t^{\lfloor n/2\rfloor})$ for $n\geq 4$, which do not scale well to high-dimensional spaces. Another approach to this problem is to approximate the convex hull by allowing a controlled level of inaccuracy Sartipizadeh and Vincent (2016); Jia et al. (2022). However, as reported in the literature Casadio et al. (2025), this method has not shown promising performance even in high-dimensional settings such as large language models, whose dimensionality is still lower than that of semantic segmentation networks.
For this reason, we keep the original formulation and retain all points in the training dataset when constructing the convex hull. Importantly, this does not create scalability issues for our experiments, since conformal inference is highly data-efficient. For example, guarantees of order $\delta_{1}=0.999$ and $\delta_{2}=0.997$ can be verified with a calibration dataset of size $m=8000$. Because we consider the size of the training dataset to be at most $50\%$ of the calibration dataset size, then $t=4000$ training samples suffice for our experiments, requiring only $4000$ coefficients $\alpha$ in the convex hull formulation. Moreover, since the number of extreme points grows exponentially with the dimension of the logits, i.e., $\mathcal{O}((\log t)^{n-1})$ Dwyer (1988), it is highly likely that most training samples will serve as extreme points. This renders attempts to distinguish between extreme and interior points largely ineffective.
[…]
In addition, we leverage conformal inference to generate a hyper-rectangle that bounds the error $q(\mathbf{vec}(x))=f(\mathbf{vec}(x))-g(\mathbf{vec}(x))$ between the models $f(\mathbf{vec}(x))$ and $g(\mathbf{vec}(x))$ with provable guarantees. This hyper-rectangle is then used to inflate the convex hull and obtain the probabilistic reachset. Here, we formulate this inflating hyper-rectangle as a convex hull:
$\mathbf{HR}=\mathcal{R}_{q}^{\epsilon}(\mathcal{W};\ \ell,m)=c+\sum_{i=1}^{n}\beta_{i}\sigma_{i}e_{i},\qquad\mathbf{s.t.}\quad\sum_{i=1}^{n}\beta_{i}=1,\quad\beta_{1},\cdots,\beta_{n}>0$
where $e_{i}\in\mathbb{R}^{n}$ is the $i$-th Cartesian unit vector in Euclidean space, $\sigma_{i}$ is the magnitude obtained from conformal inference (see Eq. (10)) along direction $e_{i}$ and $c$ is the center of the hyper-rectangle.
Given this formulation, we compute the Minkowski sum between the deterministic reach set of the surrogate model $g$ and the inflating hyper-rectangle as:
$\mathcal{R}_{f}^{\epsilon}(\mathcal{W};\ell,m)=\mathcal{R}_{g}(\mathbf{I})\oplus\mathcal{R}_{q}^{\epsilon}(\mathcal{W};\ \ell,m)=\mathbf{CH}\oplus\mathbf{HR}=c+\sum_{j=1}^{t}\alpha_{j}A\ v^{\text{train}}_{j}+\sum_{i=1}^{n}\beta_{i}\sigma_{i}e_{i},$
$\text{where,}\qquad\alpha_{1},\cdots,\alpha_{t},\beta_{1},\cdots,\beta_{n}\geq 0,\quad\sum_{j=1}^{t}\alpha_{j}=1,\quad\sum_{i=1}^{n}\beta_{i}=1.$
The projection of the probabilistic reach set $\mathcal{R}_{f}^{\epsilon}(\mathcal{W};\ell,m)$ onto each class of a given pixel can be performed efficiently. To this end, we collect the training points $v_{j}^{\text{train}},;j=1,\ldots,t$. Since the mapping from latent space to the original space is linear ($\mathbf{vec}(\hat{y})=Av$), the vertices of the convex hull in latent space are preserved under this transformation. In other words, if $v^{*}$ is an extreme point of the convex hull in latent space, then $Av^{*}$ is an extreme point of the convex hull in the original space. Therefore, we collect the points $\mathbf{vec}(\hat{y}_{j}^{\text{train}})=Av_{j}^{\text{train}}$ and compare them to determine the upper bound $\mathbf{ub}$ and lower bound $\mathbf{lb}$ of the surrogate model’s predictions. Consequently, the clipping block in the surrogate model ensures that every unseen point $\mathbf{vec}(\hat{y}^{\text{unseen}})$ satisfies
$\mathbf{lb}\leq\mathbf{vec}(\hat{y}^{\text{unseen}})<\mathbf{ub}.$
For a given pixel and class, let the corresponding logit value be the $k$-th component of $\mathbf{vec}(y^{\text{unseen}})$. Then the projection of the reachset on this component satisfies
$c(k)+\mathbf{lb}(k)-\sigma_{k}\leq\mathbf{vec}(y^{\text{unseen}})(k)\leq c(k)+\mathbf{ub}(k)+\sigma_{k}.$
We conclude this section with a comparison against our naïve approach. Specifically, we consider the experiment described in Appendix C for the given guarantee in Eq. (12), using a calibration dataset of size m=200,000m=200{,}000 and compare the resulting bounds. For this comparison, we employ a clipping block that projects with $l=\infty$. We constructed the convex hull using a training dataset of size $t=4000$. In addition, we sampled a separate dataset of size $t^{\prime}=10000$ to compute the normalization factors (see Eq. (8)) prior to performing conformal inference for obtaining the inflating hyper-rectangle. The resulting reachable set, together with the outcome of the naïve approach, in the presence of $10^{6}$ new simulations, $f(x),x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}$, are shown in Figure 4. Our calculation shows the proportion of the points lying outside of the reachset is only $0.000012$, which aligns well with the proposed $\langle\epsilon,\ell,m\rangle$ guarantee in Eq. (12).
Appendix E Additional Research Questions
Table 2: This table reports the bound ratio, $\mathsf{bound\_ratio}$, and the empirical miscoverage level of our reachset, $\hat{\epsilon}$ for a selection of test images, perturbation level and perturbation dimensions. We performed $10^{6}$ inferences to estimate the conservatism. Observing a nonzero $\hat{\epsilon}$ is a promising indication of a low level of conservatism.
Dataset Model Image name # Perturbed Pixels $e$ $\mathsf{bound\_ratio}$ $\hat{\epsilon}$
Camvid BiSeNet 0001TP_008790.png $41471\times 3$ $5/255$ 0.5328 3.08e-6
Figure 5: Shows the average run time versus perturbation level (top row) and perturbation dimension $r$ (bottom row). In the top row, the image is entirely perturbed within an $\ell_{2}$ ball for CamVid and Lung Segmentation and $\ell_{\infty}$ ball for OCTA-500. In the bottom row, we have darkening adversary where $e$ is fixed at $5/255$ across all experiments. The runtimes are averaged over 200 test images from datasets.
(a) Image, 0001TP_008790.png
(b) Predicted Mask from BiSeNet
(c) Image perturbed over $6912$ darkened pixels.
(d) Image perturbed over $41472$ darkened pixels.
(e) Unknown and non-robust masks for darkening adversary with $e=5/255$ and $r=6192\times 3$.
(f) Unknown and non-robust masks for darkening adversary with $e=5/255$ and $r=41472\times 3$.
Figure 6: Visualization for verification on BiSeNet for a test image from the CamVid dataset. (a) Test image. (b) Segmentation mask predicted by BiSeNet. (c,d) Pixels selected for perturbation (in black) on the test image (we sampled $6192$ (1%) and $41472$ (6%) pixels with R, G, B intensities above $150/255$). (e,f) Robust (white), non-robust (red), and unknown (black) pixels for perturbation magnitudes $e=5/255$ and dimensions $r=6192\times 3$ and $r=41472\times 3$.
(a) Image, CHNCXR_0005_0.png
(b) Predicted Mask from UNet1
(c) Image perturbed over $2621$ darkened pixels.
(d) Image perturbed over $15726$ darkened pixels.
(e) Unknown and non-robust masks for darkening adversary with $e=5/255$ and $r=2621\times 1$.
(f) Unknown and non-robust masks for darkening adversary with $e=5/255$ and $r=15726\times 1$.
Figure 7: Visualization for verification on UNet1 for a test image from the Lung Segmentation dataset. (a) Test image. (b) Segmentation mask predicted by UNet1. (c,d) Pixels selected for perturbation (in black) on the test image (we sampled $2621$ (1%) and $15726$ (6%) pixels with Gray intensities above $150/255$). (e,f) Robust (white), non-robust (red), and unknown (black) pixels for perturbation magnitudes $e=5/255$ and dimensions $r=2621\times 1$ and $r=15726\times 1$.
(a) Image, 10491.bmp from OCTA-500 dataset.
(b) Predicted Mask from UNet2
(c) Image perturbed over $924$ darkened pixels.
(d) Image perturbed over $5544$ darkened pixels.
(e) Unknown and non-robust masks for darkening adversary with $e=5/255$ and $r=924\times 1$.
(f) Unknown and non-robust masks for darkening adversary with $e=5/255$ and $r=5544\times 1$.
Figure 8: Visualization for verification on UNet2 for a test image from the OCTA-500 dataset. (a) Test image. (b) Segmentation mask predicted by UNet2. (c,d) Pixels selected for perturbation (in black) on the test image (we sampled $924$ (1%) and $5544$ (6%) pixels with Gray intensities above $150/255$). (e,f) Robust (white), non-robust (red), and unknown (black) pixels for perturbation magnitudes $e=5/255$ and dimensions $r=924\times 1$ and $r=5544\times 1$.
RQ2: Do State-of-the-Art Deterministic Verification Techniques Scale to the Level Achieved by Our Method on Complex and High-Dimensional SSN Models?
We applied techniques such as $\alpha$-$\beta$-CROWN (Zhou et al., 2024) and NNV (Tran et al., 2021) to our segmentation experiments using the UNet1, UNet2, and BiSeNet models in RQ1. However, due to the size of these models and the high perturbation levels considered, both methods encountered out-of-memory errors and failed to complete verification on the same hardware used for our approach. This highlights that while deterministic guarantees are often preferable, they may not always be computationally practical. In such cases, our probabilistic verification method for SSNs offers a scalable and effective alternative.
RQ3: How Does the Performance of our Methodology Compare to Existing Probabilistic Verification Methods for Neural Networks in the Literature?
In this section, we present a comparison with the works of Fiacchini and Alamo (2021); Anani et al. (2024) on the Cityscapes dataset Cordts et al. (2016). Since the formulation of verification guarantees differs slightly across these approaches, we first provide a general overview of the techniques proposed in Fiacchini and Alamo (2021); Anani et al. (2024), followed by a brief description of our own method. This will clarify how the comparison can be meaningfully established.
Overview 1. In Fischer et al. (2021), the authors introduced a probabilistic method to verify Semantic Segmentation Neural Networks. Inspired by Cohen et al. (2019), they introduced randomness via a Gaussian noise $\nu\sim\mathcal{N}(0,\sigma^{2})\in\mathbb{R}^{h\times w\times nc}$, applied to the input, and constructed a smoothed version of the segmentation model, denoted $\overline{\mathsf{SSN}}(x)(i,j)$, for each mask pixel $(i,j)$:
$\overline{\mathsf{SSN}}(x)(i,j)=c_{A}(i,j)=\underset{c\in\mathbf{L}}{\arg\max}\ \Pr_{\nu\sim\mathcal{N}(0,\sigma^{2})}\left[\mathsf{SSN}(x+\nu)(i,j)=c\right]$
They then established that, with confidence level $\delta_{2}$, the smoothed prediction $\overline{\mathsf{SSN}}(x)(i,j)$ is robust within an $\ell_{2}$ ball $\mathbf{B}_{\bar{r}_{i,j}}(x)$ of radius $\bar{r}_{i,j}=\sigma\Phi^{-1}(\underline{p_{A}}(i,j))$, where $\underline{p_{A}}(i,j)$ is a lower bound on the class probability:  $p_{A}(i,j)=\Pr_{\nu\sim\mathcal{N}(0,\sigma^{2})}\left[\mathsf{SSN}(x+\nu)(i,j)=c_{A}(i,j)\right]$. Here, $\Phi^{-1}(.)$ is the inverse CDF of normal distribution. The corresponding probabilistic guarantee becomes:
$\forall x^{\prime}\in\mathbf{B}_{\bar{r}_{i,j}}(x):\Pr\left[\overline{\mathsf{SSN}}(x^{\prime})(i,j)=c_{A}(i,j)\right]\geq\delta_{2}$
However, certifying all mask pixels simultaneously is challenging. To enable global guarantees, a more conservative definition of smoothing is adopted: a fixed threshold $\tau\in[0.5,1]$ is used for all mask locations $(i,j)$ which implies:
$\overline{\mathsf{SSN}}(x)(i,j)=c_{A}(i,j),\quad\text{where}\quad\Pr_{\nu\sim\mathcal{N}(0,\sigma^{2})}\left[\mathsf{SSN}(x+\nu)(i,j)=c_{A}(i,j)\right]\geq\tau$
This leads to a unified radius $r=\sigma\Phi^{-1}(\tau)$ for the input space. While this simplifies the formulation, it introduces the concept of abstained or uncertifiable pixels—those that cannot meet the desired confidence threshold. To mitigate the issue of multiple comparisons (i.e., union bounds), the authors propose statistical corrections such as the Bonferroni and Holm–Bonferroni methods CE (1936); Holm (1979). Assuming the set of certifiable pixels is defined as $\mathbf{CERT}=\left\{(i,j)\mid(i,j)\text{ is certifiable}\right\}$, they propose the final guarantee in the following form:
[…]
This method is referred to as $\mathsf{SEGCERTIFY}$. A key limitation of this technique is the presence of abstained pixels. The authors in Anani et al. (2024) addressed this by proposing an adaptive approach, $\mathsf{ADAPTIVECERTIFY}$, which reduces the number of uncertifiable pixels. However, the guarantees proposed in Fiacchini and Alamo (2021); Anani et al. (2024) apply only to the smoothed model, not the base model. To express this in terms of the base model, we define the following problem:
Problem 1. For a given image $x$, noise $\nu\sim\mathcal{N}(0,\sigma^{2})$, and a threshold $\tau\in[0.5,1]$, define $c_{A}(i,j)\in\mathbf{L}$ such that:  $\Pr_{\nu\sim\mathcal{N}(0,\sigma^{2})}\left[\mathsf{SSN}(x+\nu)(i,j)=c_{A}(i,j)\right]\geq\tau$. Then for any $x^{\prime}\in\mathbf{B}_{r}(x)$ with $r=\sigma\Phi^{-1}(\tau)$ and confidence level $\delta_{2}$, we want to show the following guarantee, where $\mathbf{CERT}$ will be also determined through the verification process:
$\Pr\big{[}\!\!\!\!\bigwedge_{(i,j)\in\mathbf{CERT}}\!\!\!\Pr_{\nu\sim\mathcal{N}(0,\sigma^{2})}\left[\mathsf{SSN}(x^{\prime}+\nu)(i,j)=c_{A}(i,j)\right]\geq\tau\ \ \ \ \big{]}\geq\delta_{2}$
Overview 2. In our approach, for a given image $x$ and input set $\mathbf{B}_{r}(x)$, we perform a probabilistic reachability analysis with a $\langle\epsilon,\ell,m\rangle$ guarantee. Due to the presence of conservatism in our reachability technique, some mask pixels are marked as robust (i.e., certifiable), while others cover multiple classes and are considered unknown or uncertifiable. Let $\mathbf{CONFORMAL\_CERT}=\left\{(i,j)\mid(i,j)\text{ is certifiable}\right\}$. Then our verification objective is:
Problem 2. Given an image $x$, input set $\mathbf{B}_{r}(x)$, and sampling distribution $x^{\prime}\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{B}_{r}(x)$, for coverage level $\delta_{1}=1-\epsilon\in[0,1]$, hyper-parameters $\ell,m$ and confidence level $\delta_{2}=1-\mathsf{betacdf}_{\delta_{1}}(\ell,m+1-\ell)$, we aim to show the following $\langle\epsilon,\ell,m\rangle$ guarantee, where here the set $\mathbf{CONFORMAL\_CERT}$ will be also determined through the verification process:
$\Pr\big{[}\Pr\big{[}\hskip-28.45274pt\bigwedge_{(i,j)\in\mathbf{CONFORMAL\_CERT}}\hskip-28.45274pt\mathsf{SSN}(x^{\prime})(i,j)=\mathsf{SSN}(x)(i,j)\hskip 8.53581pt\big{]}\geq\delta_{1}\big{]}\geq\delta_{2}$
Comparison. The unknown pixels in our method are conceptually equivalent to the abstained pixels in Fiacchini and Alamo (2021); Anani et al. (2024). Thus, given the same input set $\mathbf{B}_{r}(x)$ where $r$ is provided by Fiacchini and Alamo (2021), the comparison focuses to show which technique results in fewer uncertifiable mask pixels. To this end, we replicate the setup of Table 1 from Anani et al. (2024), using the same HrNetV2 model Wang et al. (2020)trained on the Cityscapes dataset with the HrNetV2-W48 backbone. In this case study, the results of our Naive approach were approximately equivalent to those of the surrogate-based approach. Therefore, we use the Naive technique for comparison, as it is more efficient than the surrogate-based method. We compare the number of uncertifiable pixels produced by our naive approach against the abstained pixels reported in their results. Our findings are summarized in Table 3, which shows the average over 200 test images. We also report our verification runtime for completeness.
Table 3: Percentage of uncertifiable pixels under different $(\sigma,\tau,r)$ settings for Fiacchini and Alamo (2021), Anani et al. (2024), and our Naive approach. The verification runtime is $210$ seconds for all experiments.
$\sigma$ $\tau$ $r$ Fiacchini and Alamo (2021) (%) Anani et al. (2024) (%) Ours (%)
0.25 0.75 0.1686 7 5 0.0642
0.33 0.75 0.2226 14 10 0.0676
0.50 0.75 0.3372 26 15 0.0705
0.25 0.95 0.4112 12 9 0.0727
0.33 0.95 0.5428 22 18 0.0732
0.50 0.95 0.8224 39 28 0.0758

### v2

@@
 Abstract
-Semantic segmentation networks (SSNs) play a critical role in domains such as medical imaging, autonomous driving, and environmental monitoring, where safety hinges on reliable model behavior under uncertainty.
-Yet, existing probabilistic verification approaches struggle to scale with the complexity and dimensionality of modern segmentation tasks, often yielding guarantees that are too conservative to be practical.
-We introduce a probabilistic verification framework that is both architecture-agnostic and scalable to high-dimensional outputs.
-Our approach combines sampling-based reachability analysis with conformal inference (CI) to deliver provable guarantees while avoiding the excessive conservatism of prior methods.
-To counteract CI’s limitations in high-dimensional settings, we propose novel strategies that reduce conservatism without compromising rigor.
-Empirical evaluation on large-scale segmentation models across CamVid, OCTA-500, Lung Segmentation, and Cityscapes demonstrates that our method provides reliable safety guarantees while substantially tightening bounds compared to SOTA.
-We also provide a toolbox implementing this technique, available on github.
-[^1] https://github.com/Navidhashemicodes/SSN_Reach_CLP_Surrogate
+Semantic segmentation networks (SSNs) are central to safety-critical applications such as medical imaging and autonomous driving, where robustness under uncertainty is essential.
+However, existing probabilistic verification methods often fail to scale with the complexity and dimensionality of modern segmentation tasks, producing guarantees that are overly conservative and of limited practical value.
+We propose a probabilistic verification framework that is architecture-agnostic and scalable to high-dimensional input-output spaces.
+Our approach employs conformal inference (CI), enhanced by a novel technique that we call the clipping block, to provide provable guarantees while mitigating the excessive conservatism of prior methods.
+Experiments on large-scale segmentation models across CamVid, OCTA-500, Lung Segmentation, and Cityscapes demonstrate that our framework delivers reliable safety guarantees while substantially reducing conservatism compared to state-of-the-art approaches on segmentation tasks.
+We also provide a public GitHub repository [^1] https://github.com/Navidhashemicodes/SSN_Reach_CLP_Surrogate for this approach, to support reproducibility.
 Keywords: Robustness, Semantic Segmentation, Verification, Conformal Inference
 1 Introduction
-Deep Neural Networks have achieved outstanding success across diverse domains, including image classification, speech recognition, medical diagnosis, and autonomous systems.
-However, despite their effectiveness, DNNs remain highly sensitive to input perturbations and adversarial examples—carefully crafted modifications that can cause significant changes in model predictions.
-This vulnerability poses serious risks, particularly in safety-critical applications such as autonomous driving, and medical imaging, where even a single misprediction may lead to catastrophic consequences.
-Research on neural network robustness has generally followed two main directions.
-Falsification-based approaches aim to uncover adversarial examples that expose a model’s vulnerabilities (Madry et al., 2017; Goodfellow et al., 2014; Kurakin et al., 2016).
-While effective for identifying potential failure cases, these methods provide no formal guarantees of safety.
-In contrast, verification-based approaches seek to certify robustness by proving that a network remains reliable under specified input perturbations (Zhou et al., 2024; Lemesle et al., 2024; Wu et al., 2024).
-Although these techniques offer provable guarantees, they are typically hindered by severe scalability challenges.
-Reachability analysis—the task of computing the set of all possible outputs of a system given a bounded input set—has become a central tool for verifying robustness.
-In the context of neural networks, it captures how input perturbations propagate through the layers and affect the final prediction.
-However, reachability analysis quickly becomes computationally infeasible for large-scale models or high-dimensional outputs.
-Semantic segmentation networks (SSNs), which assign a class label to every pixel, present a particularly difficult case due to their extremely high-dimensional output space and complex architectures, often involving components such as transposed convolutions, max-pooling, and dilated convolutions.
-For such models, traditional reachability-based methods either fail to scale or yield overly conservative guarantees.
-To overcome these challenges, probabilistic verification has emerged as a promising direction (Fischer et al., 2021; Anani et al., 2024; Hao et al., 2022; Marzari et al., 2024).
-Instead of requiring guarantees to hold for every possible perturbation, this approach establishes them with provable probability under a prescribed distribution of input uncertainty.
-Such methods tend to scale more effectively and yield informative statistical assurances, particularly when exact reachability bounds or Lipschitz estimates are impractical or overly pessimistic.
-Yet, current probabilistic techniques remain limited: they often cannot capture complex output specifications and struggle to handle the scale of modern semantic segmentation networks.
-In this paper, we introduce a scalable, architecture-independent probabilistic verification framework with reduced conservatism, specifically designed for semantic segmentation models.
-Our approach makes no restrictive assumptions on network layers or structure, and is naturally applicable to high-dimensional outputs.
+Semantic image segmentation is widely applied in high-stakes areas, including medical imaging (perone2018spinal) and autonomous driving (deng2017cnn).
+Yet, segmentation models powered by deep learning are known to be fragile, as they can be easily manipulated by adversarial inputs (xie2017adversarial), limiting their safe use in these critical applications.
+[…]
+A prominent research direction in probabilistic verification is randomized smoothing, first introduced for classifiers cohen2019certified and later adapted to segmentation models fischer2021scalable, with subsequent improvements in anani2024adaptive; hao2022gsmooth.
+More recently, conformal inference has been investigated as a means to provide probabilistic guarantees in classification tasks jeary2024verifiably.
+In hashemiscaling, the authors proposed a methodology to extend the application of conformal inference from robustness analysis in classification task to robustness analysis in segmentation tasks.
+They first introduce a naive approach that relies solely on conformal inference, which produces only hypercubes for the reachable set of the vector of logits—also described in the preliminaries of this paper.
+They then improve this method by training a small ReLU neural network as a surrogate for the base segmentation model, demonstrating that the surrogate enables the construction of more general convex shapes for the system’s reachable set.
+Appendix A presents a comprehensive comparison between segmentation verification techniques based on randomized smoothing fischer2021scalable; anani2024adaptive and the conformal inference approach in hashemiscaling, highlighting the advantages of conformal inference for segmentation tasks.
+However, as we detail later, the method in hashemiscaling has its own limitations, which we address in this paper.
+In fact, the small ReLU neural network introduced in hashemiscaling poses several challenges for this methodology.
+First, the surrogate model can suffer from fidelity issues, leading to overly conservative guarantees.
+Second, the high dimensionality of the input space makes training infeasible.
+Moreover, it imposes restrictive assumptions on the perturbation of the image, requiring it to be sparse and confined to $\ell_{\infty}$-type perturbations.
+In this paper, we propose an alternative technique to replace this ReLU surrogate model and to avoids all these issues.
+Specifically, we introduce a clipping-block at the end of the base model.
+This approach requires no surrogate training—eliminating fidelity concerns—and allows perturbations across the entire image, supporting general $\ell_{p}$ perturbations.
+To show this numerically, we deliberately consider the same models and datasets as in hashemiscaling, but with higher-dimensional (and also full) image perturbations in our numerical experiments and we observed that the surrogate-based approach in hashemiscaling fails under these conditions because deterministic reachability using the NNV toolbox tran2020nnv cannot handle the dimensionality of the input sets used in our experiments facing runtime and out-of-memory errors.
 Related Works.
-[…]
 Deterministic Verification of Neural Networks.
-Deterministic verification of neural networks aims to provide formal guarantees that the network’s outputs meet a given specification for all inputs within a prescribed set.
-This problem has been addressed using several well-established techniques, such as Satisfiability Modulo Theories (SMT) solvers (Duong et al., 2023; Katz et al., 2017; Wu et al., 2024), Mixed Integer Linear Programming (MILP) formulations (Anderson et al., 2020; Cheng et al., 2017; Tjeng et al., 2017), and different types of reachability analysis (Bak et al., 2020; Tran et al., 2020).
-Many of these methods can be viewed through the general framework of branch-and-bound algorithms (Bunel et al., 2020), which iteratively partition and prune the input domain to either uncover counterexamples or prove correctness.
-Prominent tools in this area, including $\alpha,\beta$-CROWN (Zhou et al., 2024), PyRAT (Lemesle et al., 2024), and Marabou (Wu et al., 2024), have demonstrated strong results in recent VNN-COMP competitions Brix et al.
-(2024), consistently ranking at the top of evaluation benchmarks.
-A key component of branch-and-bound approaches is computing tight output bounds for neural networks over constrained input regions.
-Techniques like interval arithmetic (Cheng et al., 2017; Pulina and Tacchella, 2010) are commonly employed to compute these bounds, influencing both the accuracy and computational efficiency of the verification.
+Deterministic verification of neural networks has been addressed using several well-established techniques, such as Satisfiability Modulo Theories (SMT) solvers (duong2023dpll; katz2017reluplex; wu2024marabou), Mixed Integer Linear Programming (MILP) formulations (anderson2020strong; cheng2017maximum; tjeng2017evaluating), and different types of reachability analysis (bak2020improved; tran2020verification).
+Many of the existing methods can be also viewed through the general framework of branch-and-bound algorithms (bunel2020branch), which iteratively partition and prune the input domain to either uncover counterexamples or prove correctness.
+Prominent tools in this area are, $\alpha,\beta$-CROWN, (zhou2024scalable), PyRAT (lemesle2024neural), and Marabou (wu2024marabou).
+Existing deterministic techniques do not scale to segmentation models because of their high level of complexity and dimensionality.
+Conformal Inference.
+The integration of CI with formal verification techniques has recently received noticeable interest, that is primarily due to its accuracy and level of scalability.
+For instance, bortolussi2019neural merges CI with neural state classifiers to develop a stochastic runtime verification algorithm.
+lindemann2023safe; zecchin2024forking employs CI to guarantee results in MPC control using a trained model.
+tonkens2023scalable applies CI for planning with probabilistic safety guarantees, and hashemi2023data; hashemi2024statistical employs CI for reachability of stochastic dynamical systems, see lindemann2024formal for a recent survey article.
 2 Preliminaries
@@
 We write $x\sim\mathcal{X}$ to indicate that the random variable $x$ is drawn from the distribution $\mathcal{X}$, and $x\stackrel{{\scriptstyle\mathcal{X}}}{{\sim}}\mathbf{X}$ to indicate that $x$ is sampled from the set $\mathbf{X}$ according to the distribution $\mathcal{X}$.
+We use $\mathsf{betacdf}$ to denote the CDF of beta distribution.
 We use bold letters to represent sets and calligraphic letters to denote distributions.
-Consider a 3-dimensional tensor $x\in\mathbb{R}^{h\times w\times nc}$, which we flatten into a vector.
+Consider a 3-dimensional tensor $x\in\mathbb{R}^{h\times w\times nc}$ of height $h$, width $w$, and $nc$ channels, which we flatten into a vector.
 The vectorized form is denoted as $\mathbf{vec}(x)\in\mathbb{R}^{n_{0}}$, where $n_{0}=nc\times w\times h$.
@@
 The ceiling operator $\lceil x\rceil$ denotes the smallest integer greater than or equal to $x\in\mathbb{R}$.
-A feedforward neural network with input dimension $n_{0}$, $l$ hidden layers of sizes $n_{1},\ldots,n_{l}$, and output dimension $N$, using the $\mathrm{ReLU}$ activation function, is denoted as $n_{0},n_{1},\ldots,n_{l},N](\mathrm{ReLU})$.
-Her we denote, a $\ell_{2}$ ball with radius $r\in\mathbb{R}^{n_{0}}$ and center $x\in\mathbb{R}^{n_{0}}$ with $\mathbf{B}_{r}(x)$.
-Semantic Segmentation Neural Networks
-A semantic segmentation neural network (SSN) is a nonlinear function that assigns a class label to each pixel $x(i,j)$ in a multichannel input image $x$, producing a target class label $\mathsf{mask}(i,j)$ from a predefined set of classes $\mathbf{L}=\left\{1,2,\ldots,L\right\}$,
-$\mathsf{SSN}:x\in\mathbb{R}^{h\times w\times nc}\mapsto\mathsf{mask}\in\mathbf{L}^{h\times w},$
-where $h$, $w$, and $nc$ denote the height, width, and number of channels of the input image, respectively, and $(i,j)\in{1,\ldots,h}\times{1,\ldots,w}$ are pixel coordinates.
-In this paper, we denote the logits of SSNs by $y\in\mathbb{R}^{h\times w\times L}$, where the target class label $\mathsf{mask}(i,j)$ is given by $\mathsf{mask}(i,j)=\underset{l\in\mathbf{L}}{\arg\max}\ \ y(i,j,l)$.
-Here, we also define the function $f:\mathbb{R}^{n_{0}}\mapsto\mathbb{R}^{n}$ as a nonlinear map between the flattened images and the flattened logits.
-Formally,
-[…]
-Conformal Inference
-Consider a collection of i.i.d and positive scalar random variables $\mathbf{M}=\left\{R_{1},R_{2},\ldots,R_{m}\right\}$ sampled from $R\sim\mathcal{D}$, where $R_{1}<R_{2}<\ldots<R_{m}$.
-Given a new draw $R_{m+1}$ from the same distribution and a miscoverage level $\epsilon\in(0,1)$, conformal inference (CI)(Vovk et al., 2005) constructs a prediction interval $C(R_{m+1})=[0,d]$ such that $\Pr[\ \ R_{m+1}\in C(R_{m+1})\ \ ]\ \geq 1-\epsilon$.
-In this paper, we refer to these scalar random variables as nonconformity scores and we refer to $\mathbf{M}$ as the calibration dataset.
-To compute the threshold $d$, we use the empirical distribution of nonconformity scores $R_{i}$ for $i\in 1,2,\ldots,m$, and define a rank $\ell:=\lceil(m+1)(1-\epsilon)\rceil$.
-A valid choice for the threshold $d$ is then $R_{\ell}$, the $\ell$-th smallest element of the augmented set $R_{1},\ldots,R_{m},\infty$}, as defined in (Tibshirani et al., 2019), which yields the following marginal guarantee:
+Here, we denote a $\ell_{2}$ ball with radius $r\in\mathbb{R}^{n_{0}}$ and center $x\in\mathbb{R}^{n_{0}}$ with $\mathbf{B}_{r}(x)$.
+Semantic Segmentation Neural Networks (SSN)
+Consider an input image $x\in\mathbb{R}^{h\times w\times nc}$.
+An SSN processes this image and outputs a $2$-D label mask, assigning a class to every pixel location $(i,j)\in\left\{1,\ldots,h\right\}\times\left\{1,\ldots,w\right\}$, we have: $\text{SSN}(x)=\text{mask}\in\mathbf{L}^{h\times w},\quad\mathbf{L}=\{1,2,\ldots,L\}$.
+The network computes the logits $y\in\mathbb{R}^{h\times w\times L}$ in the last layer.
+For each pixel, the predicted label is obtained by choosing the class with the largest logit value: $\text{mask}(i,j)=\arg\max_{\ell\in L}\;y(i,j,\ell).$
+Instead of working directly with multidimensional tensors, it is sometimes convenient to flatten both the input image and the logits.
+Let $n_{0}=h\cdot w\cdot nc$ and $n=h\cdot w\cdot L$.
+Then the map between flattened input $\mathbf{vec}(x)$ and the flattened logit values $\mathbf{vec}(y)$ can equivalently be viewed as a nonlinear function
+$f:\mathbb{R}^{n_{0}}\to\mathbb{R}^{n},\quad f(\text{vec}(x))=\text{vec}(y),$
+Conformal Inference and $\langle\epsilon,\ell,m\rangle$ guarantee
+Suppose we have $m$ i.i.d.
+non-negative random variables $\mathbf{M}=\{R_{1},R_{2},\ldots,R_{m}\},\quad R_{1}<R_{2}<\cdots<R_{m}$, drawn from some distribution $R\sim\mathcal{D}$.
+Conformal inference (CI) (vovk2005algorithmic) provides a way to construct prediction intervals with guaranteed coverage at a user-specified miscoverage level $\epsilon\in(0,1)$.
+For a new draw $R_{m+1}$ from the same distribution $R_{m+1}\sim\mathcal{D}$, CI defines a prediction interval $C(R_{m+1})=[0,d]$ such that $\Pr\big[R_{m+1}\in C(R_{m+1})\big]\geq 1-\epsilon$.
+In this setting, the values $R_{i}\in\mathbf{M}$ are often called nonconformity scores, and the set $\mathbf{M}$ is referred to as the calibration dataset.
+To compute the threshold $d$, one considers the empirical distribution of these scores.
+Specifically, define $\ell=\lceil(m+1)(1-\epsilon)\rceil$, and let $R_{\ell}$ denote the $\ell$-th smallest element ($\ell<m$) of the set $\mathbf{M}$ (tibshirani2019conformal).
+Then the prediction threshold is given by $d=R_{\ell}$, and we are guaranteed the following marginal coverage guarantee:
 $\Pr\left[R_{m+1}\leq R_{\ell}\right]\geq 1-\epsilon,$ (1)
-where the probability is over the joint randomness of both the calibration and test samples (Tibshirani et al., 2019; Vovk et al., 2005).
-The test sample $R_{m+1}$ typically refers to an unseen example drawn from the same distribution for which we aim to provide a coverage guarantee.
-In this paper, we also refer to this sample as the unseen, and denote it by $R^{\text{unseen}}$.
-Given a sampled calibration dataset, the coverage level $\delta=\Pr[R^{\text{unseen}}\leq R_{\ell}]$ is itself a random variable following a $\mathrm{Beta}(\ell,m+1-\ell)$ distribution (Angelopoulos and Bates, 2021).
-The following equations shows the mean value and the variance of the beta distribution in terms of $m,\ell$:
-$\mathbf{E}\big{[}\delta\big{]}=\frac{\ell}{m+1},\ \ \mathbf{Var}\big{[}\delta\big{]}=\frac{\ell(m+1-\ell)}{(m+1)^{2}(m+2)},$ (2)
-that shows by appropriately tuning $m$ and $\ell$, we can significantly reduce the variance $\mathbf{Var}[\delta]$, to avhieve tighter bounds.
-For instance, when $m=8000$ and $\ell=7999$, the variance of the corresponding Beta distribution is $3.123\times 10^{-8}$, which is extremely small.
-However, to present this in a completely formal way, we only considers low variance scenarios ($\mathbf{Var}(\delta)\ll 1$) and we also include this source of uncertainty by reformulating this guarantee with a double-step probablistic guarantee.
-To include this uncertainty, we can utilize the cumulative density function, CDF, of beta distribution- a function known as regularized incomplete beta function- and propose the following guarantee based on the definition of the CDF.
-$\Pr\left[\ \ \Pr[\ \ R^{\text{unseen}}\leq R_{\ell}\ \ ]>1-\epsilon\ \ \right]>1-\mathsf{betacdf}_{1-\epsilon}(\ell,m+1-\ell).$ (3)
-where $\mathsf{betacdf}_{1-\epsilon}(\ell,m+1-\ell)$ is the regularized incomplete beta function that is characterized by the choice of $m$ and $\ell$ evaluated at point $\delta_{1}=1-\epsilon$.
-The one-step marginal guarantee in (1) holds only when the rank is set to $\ell:=\lceil(m+1)(1-\epsilon)\rceil$.
-However, as shown in Appendix A, a key advantage of the double-step probabilistic guarantee is its flexibility: the strict relationship among $m$, $\ell$, and $\epsilon$ can be relaxed.
-Therefore, in this paper, we focus on the double-step guarantee and do not enforce the constraint $\ell:=\lceil(m+1)(1-\epsilon)\rceil$.
-Instead, we tune $m$, $\ell$, and $\epsilon$ such that $\mathsf{betacdf}_{1-\epsilon}(\ell,m+1-\ell)$ becomes negligibly small.
+Here, the probability is marginal which means that it is taken jointly over both the randomness of calibration dataset $\mathbf{M}$ and the test sample $R_{m+1}$ (tibshirani2019conformal; vovk2005algorithmic).
+The test sample $R_{m+1}$ is assumed to be drawn from the same underlying distribution as the calibration set.
+Following the convention in this paper, we call this new sample unseen and denote it by $R^{\text{unseen}}$.
+Since we do not assume randomness in the calibration dataset, following the proposition of vovk2005algorithmic the authors in hashemiscaling use a two-step representation of the guarantee:
+$\Pr\left[\ \ \Pr[\ \ R^{\text{unseen}}\leq R_{\ell}\ \ ]>1-\epsilon\ \ \right]>1-\mathsf{betacdf}_{1-\epsilon}(\ell,m+1-\ell),$ (2)
+where the outer probability captures the randomness in $\mathbf{M}$, while the inner probability is valid for a fixed set $\mathbf{M}$.
+For further details and clarifying examples, please see hashemiscaling.
+Here, we tune $m$, $\ell$, such that for a given $\epsilon\in[0,1]$, $\mathsf{betacdf}_{1-\epsilon}(\ell,m+1-\ell)$ becomes less than a user-specified threshold.
 Henceforth, for any logical statement $P(\ell,m)\in\left\{\mathrm{true},\mathrm{false}\right\}$, defined over the hyper-parameters $m,\ell\leq m$, we refer to the double-step guarantee,
 $\Pr\left[\ \ \Pr[\ \ P(\ell,m)\ \ ]>1-\epsilon\ \ \right]>1-\mathsf{betacdf}_{1-\epsilon}(\ell,m+1-\ell)$
-as the $\langle\epsilon,\ell,m\rangle$ guarantee and we say the statement $P(\ell,m)$ satisfies a $\langle\epsilon,\ell,m\rangle$ guarantee.
-We also refer to $\epsilon$ as the miscoverage level, $\delta_{1}:=1-\epsilon$, as the coverage level and $\delta_{2}:=1-\mathsf{betacdf}_{1-\epsilon}(\ell,m+1-\ell)$ as the confidence of guarantee.
-Appendix B provides a clarifying example on this topic.
-Determinitic and Probabilistic Reachability Analysis on Neural Networks
-Given a neural network $f:\mathbb{R}^{n_{0}}\rightarrow\mathbb{R}^{n}$ and an input set $\mathbf{I}\subset\mathbb{R}^{n_{0}}$, deterministic reachability analysis aims to construct a set $\mathbf{R}_{f}(\mathbf{I})\subset\mathbb{R}^{n}$ such that
+as the $\langle\epsilon,\ell,m\rangle$ guarantee and we say $P(\ell,m)$ satisfies a $\langle\epsilon,\ell,m\rangle$ guarantee.
+We refer to $\delta_{1}:=1-\epsilon$, as the coverage level and $\delta_{2}:=1-\mathsf{betacdf}_{1-\epsilon}(\ell,m+1-\ell)$ as the confidence level.
+Deterministic and Probabilistic Reachability Analysis on Neural Networks
+Reachability analysis for a neural network $f:\mathbb{R}^{n_{0}}\rightarrow\mathbb{R}^{n}$ has been approached in the literature in two ways.
+In deterministic reachability analysis, given an input set $\mathbf{I}\subset\mathbb{R}^{n_{0}}$, the goal is to construct a set $\mathbf{R}_{f}(\mathbf{I})\subset\mathbb{R}^{n}$ that contains all possible outputs:
 $x\in\mathbf{I}\quad\Rightarrow\quad f(x)\in\mathbf{R}_{f}(\mathbf{I}).$ (3)
-In contrast, probabilistic reachability analysis assumes a distribution $\mathcal{W}$ over the input set $\mathbf{I}$, denoted $x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}$, and for a given miscoverage level $\epsilon$, proposes a set $\mathbf{R}_{f}^{\epsilon}(\mathcal{W})\subset\mathbb{R}^{n}$ such that
+Probabilistic reachability analysis, in contrast, assumes a prior distribution $\mathcal{W}$ for inputs over $\mathbf{I}$, denoted $x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}$.
+For a given miscoverage level $\epsilon$, it aims to construct a set $\mathbf{R}_{f}^{\epsilon}(\mathcal{W})\subset\mathbb{R}^{n}$ such that the network output lies within this set with high probability:
 $x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}\quad\Rightarrow\quad\Pr\left[f(x)\in\mathbf{R}_{f}^{\epsilon}(\mathcal{W})\right]\geq 1-\epsilon.$ (4)
-In this paper, we detail how to compute a probabilistic reach set with miscoverage level $\epsilon$, and sampling distribution $x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}$ by constructing a suitable calibration set $\mathbf{M}$ of size $m$, and selecting an appropriate rank $\ell$.
-Therefore to keep the terminlogies consistent, we reformulate the probabilistic reach set in (5) using the $\langle\epsilon,\ell,m\rangle$ guarantee, and we denote it by $\mathbf{R}_{f}^{\epsilon}(\mathcal{W}\ ;\ell,m)$, which satisfies
+In this work, we detail a procedure to compute such a probabilistic reach set for a given miscoverage level $\epsilon$ and sampling distribution $x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}$.
+This involves constructing a calibration set $\mathbf{M}$ of size $m$ and selecting a rank $\ell$ to enforce the desired coverage.
+To maintain consistent terminology, we express the probabilistic reach set using the $\langle\epsilon,\ell,m\rangle$ guarantee and denote it as $\mathbf{R}_{f}^{\epsilon}(\mathcal{W}\ ;\ell,m)$, which satisfies
 $x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}\quad\Rightarrow\quad\Pr\left[\Pr\left[f(x)\in\mathbf{R}_{f}^{\epsilon}(\mathcal{W}\ ;\ell,m)\right]\geq 1-\epsilon\right]\geq 1-\mathsf{betacdf}_{1-\epsilon}(\ell,m+1-\ell).$
 Adversarial Examples and Robustness of SSNs
-An adversarial attack on an image involves perturbing the input using a set of noise images $x^{\text{noise}}_{1},\ldots,x^{\text{noise}}_{r}$ and corresponding coefficients $\lambda=[\lambda(1),\ldots,\lambda(r)]^{\top}$.
-These perturbations are applied through a parameterized function $\Delta_{\lambda,x^{\text{noise}}}(\cdot)$, generating the adversarial image as:
-$x^{\text{adv}}=\Delta_{\lambda,x^{\text{noise}}}(x)=x+\sum_{i=1}^{r}\lambda(i)x_{i}^{\text{noise}}.$ (6)
-We focus on evaluating the robustness of semantic segmentation networks (SSNs) under such adversarial settings, particularly under the unknown, bounded adversarial examples ($x^{\text{adv}}\in\mathbf{I}$) where the coefficients in vector $\lambda$ are unknown but bounded within specified ranges, i.e., $\lambda\in[\underline{\lambda},\bar{\lambda}]\subset\mathbb{R}^{r}$.
-A pixel $x(i,j)$ is deemed attacked if it faces perturbation.
-It is also considered robust if, for all adversarial examples, the predicted label remains unchanged:  $\mathsf{SSN}(\Delta_{\lambda,x^{\text{noise}}}(x))(i,j)=\mathsf{SSN}(x)(i,j).$
-Consider another scenario where there exists perturbations such that the label at $(i,j)$ changes.
-In this case, the pixel is non-robust if for all perturbations we have $\mathsf{SSN}(\Delta_{\lambda,x^{\text{noise}}}(x))(i,j)\neq\mathsf{SSN}(x)(i,j)$.
-The pixel is also unknown if for some perturbations we have $\mathsf{SSN}(\Delta_{\lambda,x^{\text{noise}}}(x))(i,j)=\mathsf{SSN}(x)(i,j)$, and for some other perturbations we have $\mathsf{SSN}(\Delta_{\lambda,x^{\text{noise}}}(x))(i,j)\neq\mathsf{SSN}(x)(i,j)$.
-To quantify robustness of SSN we consider two different metrics:
-Robustness Value.: The Robustness Value (RV) of a network is the percentage of pixels that remain robust under attack: $\quad RV=100\times(N_{\text{robust}}/N_{\text{pixels}}),\quad\text{where }N_{\text{pixels}}=h\times w$.
+Adversarial attack is a perturbation on an image by adding a combination of noise images $x^{\text{noise}}_{1},\ldots,x^{\text{noise}}_{r}$ weighted by a vector of coefficients $\lambda=[\lambda(1),\ldots,\lambda(r)]^{\top}$.
+Formally, an adversarial image is generated via
+$x^{\text{adv}}=\Delta_{\lambda,x^{\text{noise}}}(x)=x+\sum_{i=1}^{r}\lambda(i)x_{i}^{\text{noise}},\quad\lambda\in[\underline{\lambda},\bar{\lambda}]\subset\mathbb{R}^{r},$ (5)
+where the coefficients are unknown but bounded, defining the set of feasible adversarial inputs $x^{\text{adv}}\in\mathbf{I}$.
+The impact of these perturbations on semantic segmentation networks (SSNs) can be analyzed at the pixel level.
+The SSN at pixel (i,j) is robust if its predicted label is invariant under all allowed perturbations: $\mathsf{SSN}(\Delta_{\lambda,x^{\text{noise}}}(x))(i,j)=\mathsf{SSN}(x)(i,j),\quad\forall\lambda\in[\underline{\lambda},\bar{\lambda}]$.
+If the label changes under all perturbations, the pixel is non-robust: $\mathsf{SSN}(\Delta_{\lambda,x^{\text{noise}}}(x))(i,j)\neq\mathsf{SSN}(x)(i,j),\quad\forall\lambda\in[\underline{\lambda},\bar{\lambda}]$.
+A pixel is classified as unknown when some perturbations change its label while others do not.
+[…]
+Once the reachset $\mathbf{R}_{f}^{\epsilon}(\mathcal{W}\ ;\ell,m)$ over the SSN logits $y\in\mathbb{R}^{n}$ is obtained, we can analyze each pixel status by projecting it onto individual logit components.
+This projection yields, for every pixel location $(i,j)\in\{1,\ldots,h\}\times\{1,\ldots,w\}$, a set of $L$ intervals corresponding to the class labels $l\in\mathbf{L}$.
+For pixel $(i,j)$, let $l^{*}(i,j)\in\mathbf{L}$ denote the class whose logit interval has the largest lower bound.
+The pixel is classified as unknown if this lower bound does not strictly exceed the upper bounds of all other class intervals, indicating that multiple labels are possible under perturbation.
+Otherwise, if the lower bound of $l^{*}(i,j)$ is strictly greater than all others, the pixel is labeled based on the original prediction: it is robust if $l^{*}(i,j)=\mathsf{SSN}(x)(i,j)$ and non-robust if $l^{*}(i,j)\neq\mathsf{SSN}(x)(i,j)$.
+Algorithm 2 provides a detailed step-by-step procedure for computing the pixel-wise status using this approach.
 Problem Formulation
@@
 Given $K$ test images $\{x_{1},\ldots,x_{K}\}$, and a set of adversarial examples $\mathbf{I}$, considering a miscoverage level $\epsilon$, compute the average robustness value $\overline{RV}$, with strong probabilistic guarantees.
-3 Scaling Probabilistic Reachability Analysis on SSN with Strong Guarantees
-Given a specific distribution $x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}$ for sampling inputs $x$ from the input set $\mathbf{I}$, reasoning about the resulting output distribution $y\sim\mathcal{Y}$, is infeasible due to the high degree of nonlinearity in neural networks.
-As a result, providing probabilistic coverage guarantees over the network’s output space becomes a significant challenge.
-A key advantage of conformal inference is its distributional robustness—it produces valid guarantees that are robust to all family of distributions that can well represent the samples collected in the calibration dataset.
-This property makes conformal inference particularly well-suited for reachability analysis of neural networks, which are known for producing complex and often intractable output distributions.
-However, a major challenge arises when applied to high-dimensional datasets: The scalar elements $R$ in calibration dataset are defined over the outputs $y\sim\mathcal{Y}$ and the space of all possible distributions $\mathcal{Y}$ that can adequately represent the calibration dataset is significantly larger in higher dimensions and thus conformal inference tends to be overly conservative.
-This poses a significant obstacle when applying CI to semantic segmentation networks (SSNs), which are known for producing extremely high-dimensional outputs.
-In this paper, we propose a novel learning-based neural network reachability analysis method for SSNs, that leverages deflative principal component analysis (PCA) (Mackey, 2008) to address the dimensionality challenges inherent in conformal inference when applied to provide coverage guarantees on neural network with high dimensional outputs.
-In this section, we present our reachability technique in an incremental manner to clearly illustrate both the contributions and the challenges involved in applying conformal inference to SSNs.
-We begin by introducing a naive baseline approach, then highlight its limitations on high-dimensional spaces.
-Subsequently, we describe the enhancements we propose to overcome these challenges, culminating in the final version of our reachability algorithm.
-3.1 Naive Reachability Technique via Conformal Inference
-The first step to find a probabilistic reachset on neural networks with conformal inference (CI) is to generate a calibration dataset $\mathbf{M}$ of size $m$ and a training dataset $\mathbf{T}$ of size $t$.
-Generating train dataset.
-We sample $t$ different inputs $x_{j}^{\text{train}},j=1,2,\ldots,t$ from $\mathbf{I}$, with any distribution of interest, $x\stackrel{{\scriptstyle\mathcal{W^{\prime}}}}{{\sim}}\mathbf{I}$, and compute their corresponding outputs $\mathbf{vec}(y_{j}^{\text{train}})=f(\mathbf{vec}(x_{j}^{\text{train}}))$.
-This gives us the training dataset, $\mathbf{T}=\left\{(x_{1}^{\text{train}},y_{1}^{\text{train}}),(x_{2}^{\text{train}},y_{2}^{\text{train}}),\ldots,(x_{t}^{\text{train}},y_{t}^{\text{train}})\right\}$.
-Generating Calibration dataset.
-We sample the input dataset with $m$ different inputs $x^{\text{calib}}_{i},i=1,2,\ldots,m$ from the distribution $x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}$ and their corresponding set of outputs $\mathbf{vec}(y^{\text{calib}}_{i})=f(\mathbf{vec}(x^{\text{calib}}_{i})),i=1,2,\ldots,m$.
-This gives us the input/output dataset $\mathbf{IO}=\left\{(x_{1}^{\text{calib}},y^{\text{calib}}_{1}),(x_{2}^{\text{calib}},y^{\text{calib}}_{2}),\ldots,(x_{m}^{\text{calib}},y^{\text{calib}}_{m})\right\}$.
-We next attempt to use $\mathbf{IO}$ to compute a suitable nonconformity score, $R\in\mathbb{R}_{\geq 0}$.
-To that end, inspired by recent work on using conformal inference in time-series Cleaveland et al.
-(2024), we design the nonconformity scores as follows in Equation (7).
-This design provides us a hyper-rectanglular set.
-Assuming the output $\mathbf{vec}(y)=\left[\mathbf{vec}(y)(1),\ldots,\mathbf{vec}(y)(n)\right]\in\mathbb{R}^{n}$, we propose, the nonconformity scores $R^{\text{calib}}_{i}$ as,
-$R^{\text{calib}}_{i}=\max\left(\frac{|\mathbf{vec}(y^{\text{calib}}_{i})(1)-c(1)|}{\tau_{1}},\ldots,\frac{|\mathbf{vec}(y^{\text{calib}}_{i})(n)-c(n)|}{\tau_{n}}\right),\quad i=1,2,\ldots,m$ (7)
-where $c\in\mathbb{R}^{n}$ is the average of the members in the train dataset $\mathbf{T}$, and for $k=1,2,\ldots,n$, the normalization factor $\tau_{k}$ is proposed to normalize the elements of the random vectors $|\mathbf{vec}(y)-c|$ which can also be obtained using the train dataset as,
-$\tau_{k}:=\max\left(\ \ \tau^{*},\max\left(\ \ \ |\mathbf{vec}(y_{1}^{\text{train}})(k)-c(k)|,\ \ \ldots,\ \ |\mathbf{vec}(y_{t}^{\text{train}})(k)-c(k)|\ \ \ \right)\ \ \right),$ (8)
-where we set $\tau^{*}=10^{-5}\left(\sum_{k=1}^{n}\sum_{j=1}^{t}|\mathbf{vec}(y_{j}^{\text{train}})(k)-c(k)|\right)/nt$, to avoid facing zero as our normalization factor.
-We then propose the calibration dataset as $\mathbf{M}=\left\{R^{\text{calib}}_{1},R^{\text{calib}}_{2},\ldots,R^{\text{calib}}_{m}\right\}$.
-The next step is to sort the nonconformity scores in $\mathbf{M}$ based on their magnitude.
-Without loss of generality let’s assume $R^{\text{calib}}_{1}<R^{\text{calib}}_{2}<\ldots<R^{\text{calib}}_{m}$.
-Given a new draw $x^{\text{unseen}}$ from the distribution $x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}$ and the architecture of the neural network, the output $\mathbf{vec}(y^{\text{unseen}})=f(\mathbf{vec}(x^{\text{unseen}}))$ follows a specific distribution.
-We call this distribution as $y\sim\mathcal{Y}$.
-Furthermore given the map we proposed to compute the nonconformity scores, the i.i.d random variables,
-$R^{\text{unseen}}=\max\left(\frac{|\mathbf{vec}(y^{\text{unseen}})(1)-c(1)|}{\tau_{1}},\ldots,\frac{|\mathbf{vec}(y^{\text{unseen}})(n)-c(n)|}{\tau_{n}}\right)$ (9)
-adopt another distribution where we call it $R\sim\mathcal{D}$.
-Here $R^{\text{unseen}}$ and the elements $R^{\text{calib}}_{1},R^{\text{calib}}_{2},\ldots,R^{\text{calib}}_{m}$ are all sampled from the distribution $\mathcal{D}$.
-This means that we can apply conformal inference.
-Therefore, for the new draw $R^{\text{unseen}}$, given a desired rank, $\ell\leq m$ and a desired miscoverage level $\epsilon$, we can propose the following $\langle\epsilon,\ell,m\rangle$ guarantee,
-$\Pr\left[\ \ \Pr[\ \ R^{\text{unseen}}\leq R^{\text{calib}}_{\ell}\ \ ]>1-\epsilon\ \ \right]>1-\mathsf{betacdf}_{1-\epsilon}(\ell,m+1-\ell).$
-Following the definition in Equation (9), the logical statement $P_{1}(\ell,m):=[R^{\text{unseen}}\leq R_{\ell}^{\text{calib}}]$ is logically equivalent to
-$P_{2}(\ell,m):=\bigwedge_{k=1}^{n}\left[c(k)-\sigma_{k}\leq\mathbf{vec}(y^{\text{unseen}})(k)\leq c(k)+\sigma_{k}\right],\quad\text{where}\ \sigma_{k}=\tau_{k}R^{\text{calib}}_{\ell},$ (10)
-which defines a hyper-rectangular set as the reachset for any unseen output of the neural network generated by $x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\ \mathbf{I}$.
-Since $P_{1}(\ell,m)$ satisfies the $\langle\epsilon,\ell,m\rangle$ guarantee, this reachset also provides the same coverage guarantee on the unseen outputs of the neural network.
-Appendix C provides a clarifying example on this technique demonstrating that it is effective even on very deep neural networks, offering strong probabilistic guarantees with a reasonable number of required sampling.
+3 Scaling Probabilistic Reachability Analysis on SSN with Strong Guarantees Using Clipping-Block
 As the architecture of a deep neural network becomes more complex, the distribution of its output can exhibit increasingly intricate behavior.
-Conformal inference remains robust to any output distribution that represents well the calibration dataset, allowing it to handle such complexities reliably.
-However, the tightness of the reachset does not persist when the dimensionality of the network’s output increases.
+Appendix B summarizes the naive technique proposed by hashemiscaling for the reachability of SSNs.
+As it is explained in Appendix B, this technique for reachability via conformal inference remains robust to any output distribution that represents well a sampled calibration dataset, allowing it to handle such complexities reliably.
+However, the tightness of this reachset does not persist when the dimensionality of the network’s output increases.
   • The first issue is that the naive technique can only produce a hyper-rectangular reachset, that can be a significant source of conservative in high-dimensional spaces.
-  • The second issue is that the space of distributions $\mathcal{Y}$ that can provide some distribution $\mathcal{D}$ which well represent the calibration dataset is significantly larger in high dimensional space.
-3.2 From Hyper-Rectangular Reachsets to Convex Hull Representations
-Our approach to constructing a general convex hull as the reachset is based on approximating the original deep neural network $f$ with another computation graph $g:\mathbb{R}^{n_{0}}\to\mathbb{R}^{n}$, for which a guaranteed deterministic reachable set can be derived.
-The model $g$ is generated based on the training dataset $\mathbf{T}$, which implies it is accurate only for inputs $x$ sampled from the set $\mathbf{I}$.
-The core idea is to utilize its deterministic reachable set, and then inflate it to account for approximation error between $f$ and its surrogate $g$, thereby obtaining a probabilistic reachset for the original model $f$ with an $\langle\epsilon,\ell,m\rangle$ guarantee.
-Obtaining the surrogate model $g$ serves a dual purpose: it not only provides an approximation for $f$, but also defines the shape of a convex hull which when inflated serves as a tighter over-approximation of the probabilistc reachset of $f$.
-Once $g$ is computed, this convex shape can be computed as a convex hull.
-To inflate this surrogate reachset, we follow the procedure outlined in the naive technique (Section 3.1) to generate a hyper-rectangle that bounds the prediction error with a provable $\langle\epsilon,\ell,m\rangle$ guarantee.
-In other words, we define $q(\mathbf{vec}(x))=f(\mathbf{vec}(x))-g(\mathbf{vec}(x))$ and apply the naive technique on $q(\mathbf{vec}(x))$ to compute its hyper-rectangular reachset with $\langle\epsilon,\ell,m\rangle$ guarantee.
-We refer to this hyper-rectangle as inflating set.
-In conclusion, if we denote the deterministic reachset of surrogate model as the starset $\mathcal{R}_{g}(\mathbf{I})$,
+  • The second issue is that the space of distributions $\mathcal{Y}$ that can provide some distribution $\mathcal{D}$ which well represent a sampled calibration dataset is significantly larger in high dimensional space, resulting in a conservative reachable set.
+To tackle these challenges, hashemiscaling propose constructing a more flexible reachset in the form of a general convex hull, rather than a simple hyper-rectangle.
+This is achieved by approximating the original neural network $f$ with an alternative computation graph $g:\mathbb{R}^{n_{0}}\to\mathbb{R}^{n}$, for which a guaranteed deterministic reachable set can be derived.
+To obtain the surrogate model $g$ they suggest training a small ReLU neural network over the train dataset $\mathbf{T}$ and then compute its deterministic reachset over $\mathbf{I}$ using NNV toolbox tran2020nnv.
+Since $g$ is is trained on $\mathbf{T}$, we infere it is accurate only for inputs $x$ sampled from the set $\mathbf{I}$.
+The core idea is to utilize its deterministic reachable set (also known as surrogate reachset), and then inflate it to account for approximation error between $f$ and its surrogate $g$, thereby obtaining a probabilistic reachset for the original model $f$ with an $\langle\epsilon,\ell,m\rangle$ guarantee.
+To inflate the surrogate reachset, hashemiscaling adopt the procedure from the naive approach to construct a hyper-rectangle that bounds the prediction error while maintaining a provable $\langle\epsilon,\ell,m\rangle$ guarantee.
+Concretely, let $q(\mathbf{vec}(x))=f(\mathbf{vec}(x))-g(\mathbf{vec}(x))$.
+The naive technique is then applied to $q(\mathbf{vec}(x))$ to compute its hyper-rectangular reachset under the same $\langle\epsilon,\ell,m\rangle$ guarantee.
+This resulting hyper-rectangle is referred to as the inflating set.
+In conclusion, if we denote the deterministic reachset of surrogate model as $\mathcal{R}_{g}(\mathbf{I})$,
 $x\in\mathbf{I}\ \ \Rightarrow\ \ g(\mathbf{vec}(x))\in\mathcal{R}_{g}(\mathbf{I})$
@@
 $x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}\ \ \Rightarrow\ \ \Pr\left[\ \ \Pr[\ \ f(\mathbf{vec}(x))\in\mathcal{R}_{f}^{\epsilon}(\mathcal{W};\ell,m)\ \ ]>1-\epsilon\ \ \right]>1-\mathsf{betacdf}_{1-\epsilon}(\ell,m+1-\ell).$
-This technique offers two key improvements over the naive approach presented in Section 3.1:
+This technique offers two key improvements over the naive approach:
   • The reachset is no longer constrained to a hyper-rectangle, eliminating the first source of conservatism inherent in the naive method.
@@
 This implies that a useful surrogate model must maintain an acceptable level of fidelity.
-To obtain a surrogate model $g$ with high fidelity while ensuring a deterministic reachable set, we propose using the convex hull $\mathbf{CH}$ over the training outputs $y_{j}^{\text{train}},\;j=1,\dots,t$,
-$\mathbf{CH}=\sum_{j=1}^{t}\alpha_{j}\mathbf{vec}(y_{j}^{\text{train}}),\qquad\mathbf{s.t.}\quad\sum_{j=1}^{t}\alpha_{j}=1,\quad\alpha_{1},\cdots,\alpha_{t}>0$
-and then incorporating a clipping block (see Appendix D) that projects the outputs of $f$ for any unseen image $x^{\text{unseen}}$, into this convex hull.
-Denoting the clipping block with $\mathbf{CLP}$ we propose the surrogate model as: $g(\mathbf{vec}(x))=\mathbf{CLP}(f(\mathbf{vec}(x)))$, that makes certain $\mathbf{R}_{g}(\mathbf{I})=\mathbf{CH}$.
-Despite the improvements introduced in this section, reachability analysis remains unscalable in high-dimensional settings.
-The fundamental difficulty lies in the fact that projecting $\mathbf{vec}(y^{\text{unseen}})$ on $\mathbf{CH}$ does not scale well with the dimensionality of the output space.
-To mitigate this, we incorporate an scalable version of Principal Component Analysis (PCA), known as defelation algorithm, before clipping, a technique known to be effective for handling PCA in high-dimensional spaces.
-Therefore, the error $q(x)$ primarily arises from the inclusion of PCA and clipping block.
-Initialize $z_{1}\ ,\ z_{2},\ \ldots,\ z_{t}\ \leftarrow\ \mathbf{vec}(y_{1}^{\text{train}}),\ \mathbf{vec}(y_{2}^{\text{train}}),\ \ldots,\ \mathbf{vec}(y_{t}^{\text{train}})$$A\leftarrow\ \left[\ \right]$
+However, the small ReLU neural network proposed in hashemiscaling introduces three main challenges for this methodology:
+  • The surrogate model may suffer from fidelity issues, yielding correct yet conservative probabilistic reachable sets.
+  • As the perturbation dimensionality grows, the ReLU surrogate’s input dimension increases, creating scalability issues for training.
+In addition, deterministic reachability on the ReLU surrogate also scales poorly, often leading to out-of-memory errors in the NNV toolboxtran2020nnv.
+Finally, because NNV is restricted to polytopes, image perturbations must be constrained to $\ell_{\infty}$.
+In the next section, we introduce the idea of clipping-block for constructing the surrogate model $g$.
+Unlike the previous method, our new approach requires no additional training—eliminating fidelity concerns—and imposes no restrictions on the perturbation of the input image.
+3.1 Generating the Surrogate Model via Clipping-Block
+The clipping block is the core component of our surrogate model, providing both an accurate approximation of the original system and a deterministic reachable set.
+Specifically, we first generate a cloud of outputs from the training dataset and construct a convex hull enclosing these points.
+We formulate this convex hull as,
+$\mathbf{CH}=\{z\mid z=\sum_{j=1}^{t}\alpha_{j}\mathbf{vec}(y_{j}^{\text{train}}),\ \ \sum_{j=1}^{t}\alpha_{j}=1,\quad\alpha_{1},\cdots,\alpha_{t}>0\}$
+The clipping block then applies a convex optimization procedure that projects every unseen output $\mathbf{vec}(y^{\text{unseen}})$ onto this hull, as $\mathbf{vec}(\hat{y}^{\text{unseen}})=\sum_{j=1}^{t}\alpha^{*}_{j}\mathbf{vec}(y_{j}^{\text{train}})$, where,
+$\alpha^{*}_{1},\cdots,\alpha^{*}_{t}=\underset{\alpha_{1},\cdots,\alpha_{t}}{\mathbf{argmin}}\|\mathbf{vec}(y^{\text{unseen}})-\sum_{j=1}^{t}\alpha_{j}\mathbf{vec}(y_{j}^{\text{train}})\|_{l}\quad\mathbf{s.t.}\ \sum_{j=1}^{t}\alpha_{j}=1,\ \ \alpha_{1},\cdots,\alpha_{t}>0,$
+that for $l=2$ is a quadratic programming and for $l=1,\infty$ is a linear programming.
+Among the three options, we are most interested in $l=\infty$, since it minimizes the maximum error across the logit components, making it a suitable choice for reachability analysis.
+In addition, the $l=2$ formulation is far more memory-intensive compared to the $l=\infty$ case, and LP formulation scales more gracefully.
+Each problem is linear, and batching helps amortize model builds, and parallelism is also effective — which is why we observe it running stably on our setup.
+Thus, our surrogate model $g$ is composed of two main components.
+The first is the base model $f$, which takes the vectorized image $\mathbf{vec}(x^{\text{unseen}})$ and produces the corresponding logit vector $\mathbf{vec}(y^{\text{unseen}})$.
+The second is the clipping block, which takes $\mathbf{vec}(y^{\text{unseen}})$ as input and returns its projection onto the convex hull $\mathbf{CH}$, denoted as $\mathbf{vec}(\hat{y}^{\text{unseen}})$.
+In other words, if we denote the projection operation performed in the clipping block with $\mathbf{CLP}$, then,
+$g(\mathbf{vec}(x^{\text{unseen}}))=\mathbf{vec}(\hat{y}^{\text{unseen}})=\mathbf{CLP}(f(\mathbf{vec}(x^{\text{unseen}})))$
+and obviously $\mathbf{CH}$ represents the deterministic reachable set of the surrogate model, $\mathbf{R}_{g}(\mathbf{I})=\mathbf{CH}$.
+The availability of the surrogate reachset, regardless of the type or dimensionality of the input perturbation, removes the need to assume sparsity and eliminates any restrictions on the form of the image perturbation.
+Appendix C uses a toy example to make a comparison between our clipping block approach and the Naive technique on the reachability of deep neural networks.
+To generate the convex hull, it is not necessary to include the entire training dataset.
+Instead, one can identify the extreme points sartipizadeh2016computing and construct the convex hull using only those points, which reduces the number of coefficients $\alpha$ without sacrificing accuracy.
+This approach makes the convex hull more efficient to be used as it reduces the number of coefficients $\alpha$.
+Existing algorithms for identifying extreme points include Quickhull barber1996quickhull and Clarkson’s algorithm clarkson1988applications.
+Their computational complexities are $\mathcal{O}(t\log(t))$ for $n=2,3$ and $\mathcal{O}(t^{\lfloor n/2\rfloor})$ for $n\geq 4$, which do not scale well to high-dimensional spaces.
+Another approach to this problem is to approximate the convex hull by allowing a controlled level of inaccuracy sartipizadeh2016computing; jia2022investigating.
+However, as reported in the literature casadio2025nlp, this method has not shown promising performance even in high-dimensional settings such as large language models, whose dimensionality is still lower than that of semantic segmentation networks.
+For this reason, we keep the original formulation and retain all points in the training dataset when constructing the convex hull.
+Importantly, this does not create scalability issues for our experiments, since conformal inference is highly data-efficient.
+For example, guarantees of order $\delta_{1}=0.999$ and $\delta_{2}=0.997$ can be verified with a calibration dataset of size $m=8000$.
+Because we consider the size of the training dataset to be at most $50\%$ of the calibration dataset size, then $t=4000$ training samples suffice for our experiments, requiring only $4000$ coefficients $\alpha$ in the convex hull formulation.
+Moreover, since the number of extreme points grows exponentially with the dimension of the logits, i.e., $\mathcal{O}((\log t)^{n-1})$ dwyer1988average, it is highly likely that most training samples will serve as extreme points.
+This renders attempts to distinguish between extreme and interior points largely ineffective.
+Similar to hashemiscaling, which encounters scalability issues when training a surrogate model in high-dimensional output spaces, our approach also suffers from this limitation.
+Specifically, projecting $\mathbf{vec}(y^{\text{unseen}})$ onto $\mathbf{CH}$ using linear programming does not scale efficiently with output dimensionality.
+To address this, the authors of hashemiscaling introduce a scalable variant of Principal Component Analysis (PCA), known as the deflation algorithm, prior to training.
+We also found this strategy beneficial and adopt PCA before implementing the clipping block.
+Consequently, we conclude that the prediction error in our method arises primarily from the combined effects of PCA and the clipping block.
+The following section details this incorporation, which enables our technique to extend to robustness analysis of SSNs.
+Initialize $z_{1}\ ,\ z_{2},\ \ldots,\ z_{t}\ \leftarrow\ \mathbf{vec}(y_{1}^{\text{train}}),\ \mathbf{vec}(y_{2}^{\text{train}}),\ \ldots,\ \mathbf{vec}(y_{t}^{\text{train}})$ , $A\leftarrow\ \left[\ \right]$
 for $\mathsf{index}=0,1,\ldots N-1$ do
@@
 This stage aims to train the top $N$ principal directions of cloud, which are stored as columns of the matrix $A=[\vec{a}_{0},\ \vec{a}_{1},\ \ldots,\ \vec{a}_{N-1}]\in\mathbb{R}^{n\times N}$.
-We utilize this matrix to represent the high dimensional logits $y_{j}^{\text{train}}$ with a lower dimentional representative $v_{j}\in\mathbb{R}^{N}$ where $v_{j}=A^{\top}\mathbf{vec}(y_{j}^{\text{train}})$.
-Principal Component Analysis is a technique to reduce the dimensionalty, but it does not scale to high dimensional spaces.
-To address this issue, deflation algorithms have been proposed (see Algorithm 1) that compute principal components iteratively, in order of significance (Mackey, 2008).
-At each iteration, once the most dominant principal component is identified, the dataset is projected onto the orthogonal complement of that direction—effectively removing its contribution—before computing the next principal direction.
-In this paper, we adopt a learning based deflation algorithm that is presented in Algorithm 1.
-In this algorithm we set the components of the principal direction as trainable parameters and maximize the variance of the dataset in the direction of this vector.
-At each iteration of this algorithm, the optimization to train the principal direction have one global maxima one global minima and n2n-2 saddle points, which $n$ is the dimension of logits $\mathbf{vec}(y_{j}^{\text{train}})$ Mackey (2008).
-Therefore, although the training operates in a high-dimensional space, convergence to global maxima is guaranteed.
+We employ this matrix to compress the high-dimensional logits $y_{j}^{\text{train}}$ into a reduced representation $v_{j}\in\mathbb{R}^{N}$, defined as $v_{j}=A^{\top}\mathbf{vec}(y_{j}^{\text{train}})$.
+While Principal Component Analysis (PCA) is the standard tool for dimensionality reduction, its direct application does not scale effectively in very high-dimensional settings.
+To overcome this limitation, deflation-based methods have been developed (Algorithm 1).
+These algorithms extract principal components one at a time, ordered by significance (mackey2008deflation).
+After each dominant component is obtained, the dataset is projected onto its orthogonal complement, thereby eliminating its influence before proceeding to the next iteration.
+In this paper, we employ a learning-based deflation algorithm, outlined in Algorithm 1.
+In this approach, the components of the principal direction are treated as trainable parameters, and the objective is to maximize the variance of the dataset along this direction.
+At each iteration, the optimization problem for learning the principal direction admits one global maximum, one global minimum, and n2n-2 saddle points, where $n$ denotes the dimension of the logits $\mathbf{vec}(y_{j}^{\text{train}})$ mackey2008deflation.
+Consequently, despite operating in a high-dimensional space, convergence to the global maximum is guaranteed.
 Next we generate the convex hull over the cloud of point of the low dimensional representatives of logit $v_{j},j=1,\ldots,t$ and then apply the clipping block.
@@
 In conclusion we present the surrogate model for $f$ as $g(x^{\text{adv}})$ which is a scalable choice for SSNs.
-Detecting Robust, Nonrobust and Unknown Pixels with Strong Probabilistic Guarantees
-Once the reachset over the SSN logits $y\in\mathbb{R}^{n}$ is constructed as a starset, we project it onto each logit component.
-This allows us to determine the range of values for each class label $l\in\mathbf{L}$ at every pixel location $(i,j)\in\{1,\ldots,h\}\times\{1,\ldots,w\}$.
-Consequently, each pixel $(i,j)$ is associated with $L$ logit intervals, one for each class.
-Let $l^{*}\in\mathbf{L}$ denote the class whose logit interval has the highest lower bound at pixel $(i,j)$.
-If this lower bound is not strictly greater than the upper bounds of all other classes’ intervals, the pixel is labeled as unknown as it can obtain more than one labels under UBAA.
-In case it is greater, then if $l^{*}=\mathsf{SSN}(x)(i,j)$, the prediction at that pixel is labeled as robust and if $l^{*}\neq\mathsf{SSN}(x)(i,j)$, it is labeled as non-robust.
-The Algorithm 2 explains this procedure in detail.
-4 Experiments
-In our numerical evaluation, we pose three different research questions and address each of them using numerical results.
-The main one is available in this section and the rest of them are in Appendix E.
-We utilized a Linux machine, with $48$ GB of GPU memory, $512$ GB of RAM, and $112$ CPUs.
+In addition, we leverage conformal inference to generate a hyper-rectangle that bounds the error $q(\mathbf{vec}(x))=f(\mathbf{vec}(x))-g(\mathbf{vec}(x))$ between the models $f(\mathbf{vec}(x))$ and $g(\mathbf{vec}(x))$ with provable guarantees.
+This hyper-rectangle is then used to inflate the convex hull and obtain the probabilistic reachset.
+Here, we formulate this inflating hyper-rectangle as a convex hull:
+$\mathbf{HR}=\mathcal{R}_{q}^{\epsilon}(\mathcal{W};\ \ell,m)=\{z\mid z=c+\sum_{i=1}^{n}\beta_{i}\sigma_{i}e_{i},\qquad\sum_{i=1}^{n}\beta_{i}=1,\quad\beta_{1},\cdots,\beta_{n}>0\}$
+where $e_{i}\in\mathbb{R}^{n}$ is the $i$-th Cartesian unit vector in Euclidean space, $\sigma_{i}$ is the magnitude obtained from conformal inference (see Eq.
+(9)) along direction $e_{i}$ and $c$ is the center of the hyper-rectangle.
+Given this formulation, we compute the Minkowski sum between the deterministic reach set of the surrogate model $g$ and the inflating hyper-rectangle as:
+$\mathcal{R}_{f}^{\epsilon}(\mathcal{W};\ell,m)=\mathcal{R}_{g}(\mathbf{I})\oplus\mathcal{R}_{q}^{\epsilon}(\mathcal{W};\ \ell,m)=\mathbf{CH}\oplus\mathbf{HR}=$
+$\{z\mid z=c+\sum_{j=1}^{t}\alpha_{j}A\ v^{\text{train}}_{j}+\sum_{i=1}^{n}\beta_{i}\sigma_{i}e_{i},\ \ \alpha_{1},\cdots,\alpha_{t},\beta_{1},\cdots,\beta_{n}\geq 0,\ \sum_{j=1}^{t}\alpha_{j}=1,\ \sum_{i=1}^{n}\beta_{i}=1\}$
+The projection of the probabilistic reach set $\mathcal{R}_{f}^{\epsilon}(\mathcal{W};\ell,m)$ onto each class of a given pixel can be performed efficiently.
+To this end, we collect the training points $v_{j}^{\text{train}},;j=1,\ldots,t$.
+Since the mapping from latent space to the original space is linear ($\mathbf{vec}(\hat{y})=Av$), the vertices of the convex hull in latent space are preserved under this transformation.
+In other words, if $v^{*}$ is an extreme point of the convex hull in latent space, then $Av^{*}$ is an extreme point of the convex hull in the original space.
+Therefore, we collect the points $\mathbf{vec}(\hat{y}_{j}^{\text{train}})=Av_{j}^{\text{train}}$ and compare them to determine the upper bound $\mathbf{ub}$ and lower bound $\mathbf{lb}$ of the surrogate model’s predictions.
+Consequently, the clipping block in the surrogate model ensures that every unseen point $\mathbf{vec}(\hat{y}^{\text{unseen}})$ satisfies, $\mathbf{vec}(\hat{y}^{\text{unseen}})\in[\mathbf{lb},\mathbf{ub}]$.
+For a given pixel and class, let the corresponding logit value be the $k$-th component of $\mathbf{vec}(y^{\text{unseen}})$.
+Then the projection of the reachset on this component satisfies: $\mathbf{vec}(y^{\text{unseen}})(k)\in[c(k)+\mathbf{lb}(k)-\sigma_{k}\ \ ,\ \ c(k)+\mathbf{ub}(k)+\sigma_{k}]$.
+We stress that setting $\mathcal{W}^{\prime}=\mathcal{W}$ and $t=m/2$ markedly improves the surrogate model’s fidelity in our new approach a setting we always follow in our experiments via clipping block technique.
 Input: $\mathbf{I}$, $x$, $f$, $\mathcal{W}$, $\epsilon$, Hyper-parameters(,m)($\ell$,m)
 Output: The status of pixels
-// $\mathtt{Project\ the\ Starset}$
+1mm // $\mathtt{Project\ the\ reachset}$
 $\underline{y},\overline{y}]\leftarrow\mathcal{R}_{f}^{\epsilon}(\mathcal{W};\ell,m)$
+1mm
 foreach $(i,j)\in\{1{:}h\}\times\{1{:}w\}$ do
@@
 The lower bound $\mathbf{LB}$ represents the maximum darkening (channel intensities set to zero), while the upper bound $\mathbf{UB}$ corresponds to minimum darkening.
+4 Experiments
+In our numerical evaluation, we pose two different research questions and address each of them using numerical results.
+The main one is available in this section and the rest of them are in Appendix D.
+We utilized a Linux machine, with $48$ GB of GPU memory, $512$ GB of RAM, and $112$ CPUs.
 Figure 2: Shows $\overline{\mathbf{RV}}$ versus perturbation level and perturbation dimension.
@@
 One important point is, our verification runtime depends mainly on inference time, and hyperparameters ($\ell,m$), and is only slightly sensitive to perturbation magnitude and dimension $(e,r)$.
-For per-experiment runtimes, see Figure 5 in Appendix E.
+For per-experiment runtimes, see Figure 4 in Appendix D.
 Dataset Model Input Output Number of $(m,m-\ell)$ coverage confidence Average
@@
 CamVid BiSeNet $720\!\times\!960\!\times\!3$ $720\!\times\!960\!\times\!12$ $12,511,084$ $(8\mathbf{e}3,1)$ 0.999 0.997 18.3 min (Surrogate)
-Cityscapes (Appendix E) HRNetV2 $1024\!\times\!2048\!\times\!3$ $256\!\times\!512\!\times\!19$ $65,859,379$ $(921,1)$ 0.99 0.997 3.5 min (Naive)
-RQ1: How does the technique scale with model size, number of perturbed pixels, and perturbation level?
-How does output high dimensionality, and the irregularity in output distribution of large and complex NNs make our guarantees overly conservative?
-To assess scalability, we test our verification method on high-dimensional datasets (e.g., Lung Segmentation (Jaeger et al., 2014; Candemir et al., 2014), OCTA-500 (Li et al., 2019), CamVid (Brostow et al., 2009)) using large pre-trained models.
+Cityscapes (Appendix D) HRNetV2 $1024\!\times\!2048\!\times\!3$ $256\!\times\!512\!\times\!19$ $65,859,379$ $(921,1)$ 0.99 0.997 3.5 min (Naive)
+RQ1: How does the technique scale with model size and dimensionality, number of perturbed pixels, and perturbation level?
+To assess scalability, we test our verification method on high-dimensional datasets (e.g., Lung Segmentation (Jaeger2014lungdata1; Candemir2014lungdata2), OCTA-500 (j.media.2024.103092-19), CamVid (Brostow2009camvid)) using large pre-trained models.
 We evaluate performance on a subset of $200$ test images across varying perturbation dimensions and magnitudes, with results shown in Figure 2.
-Model details and probabilistic guarantees are summarized in Table 1, where we show the following $\langle\epsilon,\ell,m\rangle$ guarantee:
+Model details and probabilistic guarantees are summarized in Table 1, where we show the following $\langle\epsilon,\ell,m\rangle$ guarantee [^3] The surrogate-based technique in hashemiscaling could not handle our perturbation dimensions.:
 […]
@@
 This bound defines the perturbation space $x^{\text{adv}}\in\mathbf{I}$.
-We also present a demo for the status of all pixels in the segmentation mask in Figures 6,7 and 8.
+We also present a demo for the status of all pixels in the segmentation mask in Figures 5,6 and 7.
 To assess the conservatism, we examine the projection bounds $\underline{y},\overline{y}]$ introduced in Algorithm 2.
@@
 The degree of conservatism can be assessed by comparing $\underline{\hat{y}},\overline{\hat{y}}$] with $\underline{y},\overline{y}]$.
-To measure this, we compute $\mathsf{bound\_ratio}=\sum_{k=1}^{h\times w\times L}(\overline{\hat{y}}(k)-\underline{\hat{y}}(k))/\sum_{k=1}^{h\times w\times L}(\overline{y}(k)-\underline{y}(k))$ and due to the high cost of $10^{6}$ inferences on the models, we only perform the conservatism analysis on one specific case with BiSeNet from Figure 2, Table 2 in Appendix E details the numerical results.
+To measure this, we compute $\mathsf{bound\_ratio}=\sum_{k=1}^{h\times w\times L}(\overline{\hat{y}}(k)-\underline{\hat{y}}(k))/\sum_{k=1}^{h\times w\times L}(\overline{y}(k)-\underline{y}(k))$ and due to the high cost of $10^{6}$ inferences on the models,we only perform the conservatism analysis on one specific case with BiSeNet from Figure 2, Table 3 in Appendix D details the numerical results.
 5 Conclusion
-Our probabilistic verification method is both scalable and data-efficient.
-While it does not offer deterministic guarantees, it can provide meaningful assurances in regions where existing deterministic tools are not applicable due to computational challenges.
-In this paper, we also demonstrated this capability through a series of numerical experiments.
-Appendix A Double-Step Probabilistic Guarantee
-The necessary steps to establish the two-sided probabilistic guarantee, along with the flexibility in choosing the dataset size $m$, rank $\ell$, and failure probability $\epsilon$, begin with the theory of order statistics David and Nagaraja (2004).
-This flexibility means that the double-step probability constraint remains valid regardless of the particular values chosen for $m$, $\ell\leq m$, and $\epsilon\in[0,1]$.
-The following lemma introduces the concept of order statistics.
-Lemma 1 (From David and Nagaraja (2004))
-Consider $m$ independent and identically distributed (i.i.d.), real-valued data points $x\in[0,1]$ drawn from uniform distribution $x\sim\mathcal{U}$.
-Suppose we sort them in ascending order and denote the $\ell^{th}$ smallest number by $x_{\ell}$,(i.e., we have $x_{1}<x_{2}<\ldots<x_{m}$).
-Let $\mathbf{Beta}(\alpha,\beta)$ denote the Beta distribution [^3] The Beta distribution is a family of continuous probability distributions defined on the interval $0\leq x\leq 1$ with shape parameters $\alpha$ and $\beta$, and with probability density function $f(x;\alpha,\beta)=\frac{x^{\alpha-1}(1-x)^{\beta-1}}{B(\alpha,\beta)}$, where the constant $B(\alpha,\beta)=\frac{\Gamma(\alpha)\Gamma(\beta)}{\Gamma(\alpha+\beta)}$ and $\Gamma(z)=\int_{0}^{\infty}t^{z-1}e^{-t}\mathbf{d}t$ is the Gamma function..
-Then the uniform random variable at rank $\ell$ follows:
-$x_{\ell}\sim\mathbf{Beta}(\ell,m+1-\ell)$
-In the context of conformal inference Lemma 1, has been extended from a uniform and domain-bounded distribution to a general distribution as follows:
-Lemma 2 (From Vovk et al. (2005))
-Consider $m$ independent and identically distributed (i.i.d.), real-valued data points $R$ drawn from some distribution $R\sim\mathcal{D}$.
-Suppose we sort them in ascending order and denote the $\ell^{th}$ smallest number by $R_{\ell}$, (i.e., we have $R_{1}<R_{2}<\ldots<R_{m}$).
-Let $\mathbf{Beta}(\alpha,\beta)$ denote the Beta distribution.
-Then, for an arbitrary $R_{m+1}$ drawn from the same distribution $\mathcal{D}$, the following holds:
-$\Pr\left[R_{m+1}<R_{\ell}\right]\sim\mathbf{Beta}(\ell,m+1-\ell),\quad 1\leq\ell\leq m.$ (11)
-This result is motivated by two key observations: (1) for a continuous distribution $R\sim\mathcal{D}$, the cumulative probability $\Pr[R\leq r]$ is uniformly distributed on $[0,1]$, and (2) sorting the samples $R_{1}<R_{2}<\ldots<R_{m}$ implies that the corresponding cumulative probabilities $\Pr[R\leq R_{\ell}]$ are also ordered, i.e., $\Pr[R\leq R_{1}]<\Pr[R\leq R_{2}]<\ldots<\Pr[R\leq R_{m}]$.
-These observations implies that for all $\ell=0,1,2,\ldots,m$, the random variable $\delta:=\Pr[R\leq R_{\ell}]$ satisfies the conditions outlined in Lemma 1.
-In simple terms, this means that while a fixed dataset $R_{1},R_{2},\ldots,R_{m}$ is not sufficient to determine the exact value of $\delta=\Pr[R<R_{\ell}]$, we can infer that the value of $\delta$ must follow a Beta distribution that is solely defined based on $m$ and $\ell$.
-This allows us to tune the parameters $m$ and $\ell$ to provide strong guarantees for the coverage $R<R_{\ell}$ where $R_{\ell}$ is obtained from our fixed sampled dataset.
-This guarantee can be equivalently expressed as a two-sided probabilistic bound.
-Specifically, let $\mathsf{betacdf}_{\delta}(\ell,m+1-\ell)$ denote the cumulative distribution function (CDF) of the Beta distribution—also known as the regularized incomplete beta function—evaluated at $\delta$.
-Then, based on the definition of cumulative density function, CDF, for any confidence $\delta_{1}\in[0,1]$ and any arbitrary distribution $R\sim\mathcal{D}$, the following inequality holds:
-$\Pr\left[\ \ \Pr[\ \ R\leq R_{\ell}\ \ ]>\delta_{1}\ \ \right]>1-\mathsf{betacdf}_{\delta_{1}}(\ell,m+1-\ell).$
-Note that the key assumption is that all variables $R,R_{1},R_{2},\ldots,R_{m}$ are independently drawn from the same distribution $\mathcal{D}$, and that the previously sampled values are sorted such that $R_{1}<R_{2}<\ldots<R_{m}$.
-Importantly, the resulting probabilistic guarantee remains valid for any arbitrary choice of sample size $m$, rank $\ell$, and confidence level $\delta_{1}$, provided they lie within appropriate bounds.
-Appendix B A Clarifying Example for Conformal Inference
-Assume an arbitrary distribution $R\sim\mathcal{D}$.
-We sample a calibration dataset of size $m=100,000$, we sort the samples as $R_{1}<R_{2}<\ldots<R_{100,000}$ and we select the sample at rank $\ell=99,999$.
-In this case, given an unseen draw $R^{\text{unseen}}$ from $\mathcal{D}$ for a miscoverage level $\epsilon=0.0001$, the $\langle\epsilon,\ell,m\rangle$ guarantee for logical statement $P(\ell,m)=[R^{\text{unseen}}<R_{\ell}]$ is:
-$\Pr\left[\ \ \Pr[\ \ R^{\text{unseen}}<R_{99,999}\ \ ]\ >\ 0.9999\ \ \right]>\ 0.9995008$
-which is valid for any calibration dataset $\mathbf{M}=\left\{R_{1},R_{2},\ldots,R_{100,000}\right\}$ sampled from distribution $\mathcal{D}$ if and only if $R^{\text{unseen}}$ is also sampled from $\mathcal{D}$, and $R_{99,999}$ is the $99,999^{\text{th}}$ smallest member of $\mathbf{M}$.
-Appendix C A Clarifying Example on Naive Approach
-Figure 3: The figure depicts two reachable sets corresponding to different $\langle\epsilon,\ell,m\rangle$ guarantees.
-Both $S_{1}$ and $S_{2}$ are obtained using the naïve approach, with $S_{2}$ providing a stronger guarantee than $S_{1}$.
-Figure 4: The figure compares two reachable sets with the same $\langle\epsilon,\ell,m\rangle$ guarantees.
-$S_{1}$ is obtained using the naïve approach, while $S_{3}$ is produced by the surrogate-based technique, illustrating the accuracy of the latter.
-Since the naive technique is architecture-agnostic we apply that for the reachability analysis of a feedforward $\mathrm{ReLU}$ neural network with $60$ hidden-layers of width $100$, an input layer of dimension $784$ and an output layer of dimension $2$.
-The set of inputs is $\mathbf{I}:=[0,1]^{784}$ and we generate both calibration and train datasets by sampling $\mathbf{I}$ uniformly.
-e.g., $\mathcal{W},\mathcal{W}^{\prime}$ are both uniform distributions.
-1:  We firstly generate a calibration dataset of size $m=200,000$ and a train dataset of size $t=10,000$, we also consider the rank $\ell=199,998$ and we target the miscoverage level of $\epsilon=0.0001$.
-In this case we compute a reachable set, $\mathbf{S}_{1}$ with the following $\langle\epsilon,\ell,m\rangle$ guarantee.
-$x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}\ \ \Rightarrow\ \ \Pr\left[\ \ \Pr[\ \ f(x)\in\mathbf{S}_{1}\ \ ]>0.9999\ \ \right]>0.9999995.$ (12)
-2:  Secondly, we set $m=8464287$, $\ell=8464286$ and $\epsilon=0.000002$ which results in another reachset $\mathbf{S}_{2}$ with the following $\langle\epsilon,\ell,m\rangle$ guarantee.
-$x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}\ \ \Rightarrow\ \ \Pr\left[\ \ \Pr[\ \ f(x)\in\mathbf{S}_{2}\ \ ]>0.999998\ \ \right]>0.9999992.$ (13)
-Figure 4 shows the reachable sets $\mathbf{S}_{1}$ and $\mathbf{S}_{2}$ in the presence of $10^{6}$ new simulations, $f(x),x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}$, for validation.
-Our calculations show that the proportion of points lying outside of $\mathbf{S}_{1}$ and $\mathbf{S}_{2}$ are 0.000013 and 0.000001, respectively, which aligns well with the proposed $\langle\epsilon,\ell,m\rangle$ guarantees.
-Remark 1
-One might argue that the dimensionality issue can be addressed by partitioning the output layer and performing reachability analysis on each partition separately.
-However, this approach is suboptimal, as the overall probabilistic guarantee deteriorates.
-Specifically, the combined guarantee is governed by a union bound over the individual partitions, which can lead to an unacceptably loose $\langle\epsilon,\ell,m\rangle$ guarantee.
-Appendix D More Details on Clipping Block, Surrogate Model & Minkowski Sum
-The clipping block is the core component of our surrogate model, providing both an accurate approximation of the original system and a deterministic reachable set.
-Specifically, we first generate a cloud of outputs from the training dataset and construct a convex hull enclosing these points.
-We formulate this convex hull as,
-$\mathbf{CH}=\sum_{j=1}^{t}\alpha_{j}\mathbf{vec}(y_{j}^{\text{train}}),\qquad\mathbf{s.t.}\quad\sum_{j=1}^{t}\alpha_{j}=1,\quad\alpha_{1},\cdots,\alpha_{t}>0$
-The clipping block then applies a convex optimization procedure that projects every unseen output $\mathbf{vec}(y^{\text{unseen}})$ onto this hull, as $\mathbf{vec}(\hat{y}^{\text{unseen}})=\sum_{j=1}^{t}\alpha^{*}_{j}\mathbf{vec}(y_{j}^{\text{train}})$, where,
-$\alpha^{*}_{1},\cdots,\alpha^{*}_{t}=\underset{\alpha_{1},\cdots,\alpha_{t}}{\mathbf{argmin}}\|\mathbf{vec}(y^{\text{unseen}})-\sum_{j=1}^{t}\alpha_{j}\mathbf{vec}(y_{j}^{\text{train}})\|_{l}\quad\mathbf{s.t.}\ \sum_{j=1}^{t}\alpha_{j}=1,\ \ \alpha_{1},\cdots,\alpha_{t}>0,$
-that for $l=2$ is a quadratic programming and for $l=1,\infty$ is a linear programming.
-Among the three options, we are most interested in $l=\infty$, since it minimizes the maximum error across the logit components, making it a suitable choice for reachability analysis.
-In addition, the $l=2$ formulation is far more memory-intensive compared to the $l=\infty$ case.
-When we ran it on the OCTA-500 dataset using the Vortex machine (540 GB RAM), it resulted in an out-of-memory error.
-We also attempted the PGD variant of quadratic programming for $l=2$; however, this led to severe CPU thread oversubscription issues.
-Thus, the LP formulation scales more gracefully.
-Each problem is linear, and batching helps amortize model builds, and parallelism is also effective — which is why we observe it running stably on our setup.
-Thus, our surrogate model $g$ is composed of two main components.
-The first is the base model $f$, which takes the vectorized image $\mathbf{vec}(x^{\text{unseen}})$ and produces the corresponding logit vector $\mathbf{vec}(y^{\text{unseen}})$.
-The second is the clipping block, which takes $\mathbf{vec}(y^{\text{unseen}})$ as input and returns its projection onto the convex hull $\mathbf{CH}$, denoted as $\mathbf{vec}(\hat{y}^{\text{unseen}})$.
-In other words,
-$g(\mathbf{vec}(x^{\text{unseen}}))=\mathbf{vec}(\hat{y}^{\text{unseen}})$
-and obviously $\mathbf{CH}$ represents the deterministic reachable set of the surrogate model, $\mathbf{R}_{g}(\mathbf{I})=\mathbf{CH}.$
-To generate the convex hull, it is not necessary to include the entire training dataset.
-Instead, one can identify the extreme points Sartipizadeh and Vincent (2016) and construct the convex hull using only those points, which reduces the number of coefficients $\alpha$ without sacrificing accuracy.
-This approach makes the convex hull more efficient to be used as it reduces the number of coefficients $\alpha$.
-Existing algorithms for identifying extreme points include Quickhull Barber et al.
-(1996) and Clarkson’s algorithm Clarkson (1988).
-Their computational complexities are $\mathcal{O}(t\log(t))$ for $n=2,3$ and $\mathcal{O}(t^{\lfloor n/2\rfloor})$ for $n\geq 4$, which do not scale well to high-dimensional spaces.
-Another approach to this problem is to approximate the convex hull by allowing a controlled level of inaccuracy Sartipizadeh and Vincent (2016); Jia et al.
-(2022).
-However, as reported in the literature Casadio et al.
-(2025), this method has not shown promising performance even in high-dimensional settings such as large language models, whose dimensionality is still lower than that of semantic segmentation networks.
-For this reason, we keep the original formulation and retain all points in the training dataset when constructing the convex hull.
-Importantly, this does not create scalability issues for our experiments, since conformal inference is highly data-efficient.
-For example, guarantees of order $\delta_{1}=0.999$ and $\delta_{2}=0.997$ can be verified with a calibration dataset of size $m=8000$.
-Because we consider the size of the training dataset to be at most $50\%$ of the calibration dataset size, then $t=4000$ training samples suffice for our experiments, requiring only $4000$ coefficients $\alpha$ in the convex hull formulation.
-Moreover, since the number of extreme points grows exponentially with the dimension of the logits, i.e., $\mathcal{O}((\log t)^{n-1})$ Dwyer (1988), it is highly likely that most training samples will serve as extreme points.
-This renders attempts to distinguish between extreme and interior points largely ineffective.
-[…]
-In addition, we leverage conformal inference to generate a hyper-rectangle that bounds the error $q(\mathbf{vec}(x))=f(\mathbf{vec}(x))-g(\mathbf{vec}(x))$ between the models $f(\mathbf{vec}(x))$ and $g(\mathbf{vec}(x))$ with provable guarantees.
-This hyper-rectangle is then used to inflate the convex hull and obtain the probabilistic reachset.
-Here, we formulate this inflating hyper-rectangle as a convex hull:
-$\mathbf{HR}=\mathcal{R}_{q}^{\epsilon}(\mathcal{W};\ \ell,m)=c+\sum_{i=1}^{n}\beta_{i}\sigma_{i}e_{i},\qquad\mathbf{s.t.}\quad\sum_{i=1}^{n}\beta_{i}=1,\quad\beta_{1},\cdots,\beta_{n}>0$
-where $e_{i}\in\mathbb{R}^{n}$ is the $i$-th Cartesian unit vector in Euclidean space, $\sigma_{i}$ is the magnitude obtained from conformal inference (see Eq.
-(10)) along direction $e_{i}$ and $c$ is the center of the hyper-rectangle.
-Given this formulation, we compute the Minkowski sum between the deterministic reach set of the surrogate model $g$ and the inflating hyper-rectangle as:
-$\mathcal{R}_{f}^{\epsilon}(\mathcal{W};\ell,m)=\mathcal{R}_{g}(\mathbf{I})\oplus\mathcal{R}_{q}^{\epsilon}(\mathcal{W};\ \ell,m)=\mathbf{CH}\oplus\mathbf{HR}=c+\sum_{j=1}^{t}\alpha_{j}A\ v^{\text{train}}_{j}+\sum_{i=1}^{n}\beta_{i}\sigma_{i}e_{i},$
-$\text{where,}\qquad\alpha_{1},\cdots,\alpha_{t},\beta_{1},\cdots,\beta_{n}\geq 0,\quad\sum_{j=1}^{t}\alpha_{j}=1,\quad\sum_{i=1}^{n}\beta_{i}=1.$
-The projection of the probabilistic reach set $\mathcal{R}_{f}^{\epsilon}(\mathcal{W};\ell,m)$ onto each class of a given pixel can be performed efficiently.
-To this end, we collect the training points $v_{j}^{\text{train}},;j=1,\ldots,t$.
-Since the mapping from latent space to the original space is linear ($\mathbf{vec}(\hat{y})=Av$), the vertices of the convex hull in latent space are preserved under this transformation.
-In other words, if $v^{*}$ is an extreme point of the convex hull in latent space, then $Av^{*}$ is an extreme point of the convex hull in the original space.
-Therefore, we collect the points $\mathbf{vec}(\hat{y}_{j}^{\text{train}})=Av_{j}^{\text{train}}$ and compare them to determine the upper bound $\mathbf{ub}$ and lower bound $\mathbf{lb}$ of the surrogate model’s predictions.
-Consequently, the clipping block in the surrogate model ensures that every unseen point $\mathbf{vec}(\hat{y}^{\text{unseen}})$ satisfies
-$\mathbf{lb}\leq\mathbf{vec}(\hat{y}^{\text{unseen}})<\mathbf{ub}.$
-For a given pixel and class, let the corresponding logit value be the $k$-th component of $\mathbf{vec}(y^{\text{unseen}})$. Then the projection of the reachset on this component satisfies
-$c(k)+\mathbf{lb}(k)-\sigma_{k}\leq\mathbf{vec}(y^{\text{unseen}})(k)\leq c(k)+\mathbf{ub}(k)+\sigma_{k}.$
-We conclude this section with a comparison against our naïve approach.
-Specifically, we consider the experiment described in Appendix C for the given guarantee in Eq.
-(12), using a calibration dataset of size m=200,000m=200{,}000 and compare the resulting bounds.
-For this comparison, we employ a clipping block that projects with $l=\infty$.
-We constructed the convex hull using a training dataset of size $t=4000$.
-In addition, we sampled a separate dataset of size $t^{\prime}=10000$ to compute the normalization factors (see Eq.
-(8)) prior to performing conformal inference for obtaining the inflating hyper-rectangle.
-The resulting reachable set, together with the outcome of the naïve approach, in the presence of $10^{6}$ new simulations, $f(x),x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}$, are shown in Figure 4.
-Our calculation shows the proportion of the points lying outside of the reachset is only $0.000012$, which aligns well with the proposed $\langle\epsilon,\ell,m\rangle$ guarantee in Eq.
-(12).
-Appendix E Additional Research Questions
-Table 2: This table reports the bound ratio, $\mathsf{bound\_ratio}$, and the empirical miscoverage level of our reachset, $\hat{\epsilon}$ for a selection of test images, perturbation level and perturbation dimensions.
-We performed $10^{6}$ inferences to estimate the conservatism.
-Observing a nonzero $\hat{\epsilon}$ is a promising indication of a low level of conservatism.
-Dataset Model Image name # Perturbed Pixels $e$ $\mathsf{bound\_ratio}$ $\hat{\epsilon}$
-Camvid BiSeNet 0001TP_008790.png $41471\times 3$ $5/255$ 0.5328 3.08e-6
-Figure 5: Shows the average run time versus perturbation level (top row) and perturbation dimension $r$ (bottom row).
-In the top row, the image is entirely perturbed within an $\ell_{2}$ ball for CamVid and Lung Segmentation and $\ell_{\infty}$ ball for OCTA-500.
-In the bottom row, we have darkening adversary where $e$ is fixed at $5/255$ across all experiments.
-The runtimes are averaged over 200 test images from datasets.
-(a) Image, 0001TP_008790.png
-(b) Predicted Mask from BiSeNet
-(c) Image perturbed over $6912$ darkened pixels.
-(d) Image perturbed over $41472$ darkened pixels.
-(e) Unknown and non-robust masks for darkening adversary with $e=5/255$ and $r=6192\times 3$.
-(f) Unknown and non-robust masks for darkening adversary with $e=5/255$ and $r=41472\times 3$.
-Figure 6: Visualization for verification on BiSeNet for a test image from the CamVid dataset.
-(a) Test image.
-(b) Segmentation mask predicted by BiSeNet.
-(c,d) Pixels selected for perturbation (in black) on the test image (we sampled $6192$ (1%) and $41472$ (6%) pixels with R, G, B intensities above $150/255$).
-(e,f) Robust (white), non-robust (red), and unknown (black) pixels for perturbation magnitudes $e=5/255$ and dimensions $r=6192\times 3$ and $r=41472\times 3$.
-(a) Image, CHNCXR_0005_0.png
-(b) Predicted Mask from UNet1
-(c) Image perturbed over $2621$ darkened pixels.
-(d) Image perturbed over $15726$ darkened pixels.
-(e) Unknown and non-robust masks for darkening adversary with $e=5/255$ and $r=2621\times 1$.
-(f) Unknown and non-robust masks for darkening adversary with $e=5/255$ and $r=15726\times 1$.
-Figure 7: Visualization for verification on UNet1 for a test image from the Lung Segmentation dataset.
-(a) Test image.
-(b) Segmentation mask predicted by UNet1.
-(c,d) Pixels selected for perturbation (in black) on the test image (we sampled $2621$ (1%) and $15726$ (6%) pixels with Gray intensities above $150/255$).
-(e,f) Robust (white), non-robust (red), and unknown (black) pixels for perturbation magnitudes $e=5/255$ and dimensions $r=2621\times 1$ and $r=15726\times 1$.
-(a) Image, 10491.bmp from OCTA-500 dataset.
-(b) Predicted Mask from UNet2
-(c) Image perturbed over $924$ darkened pixels.
-(d) Image perturbed over $5544$ darkened pixels.
-(e) Unknown and non-robust masks for darkening adversary with $e=5/255$ and $r=924\times 1$.
-(f) Unknown and non-robust masks for darkening adversary with $e=5/255$ and $r=5544\times 1$.
-Figure 8: Visualization for verification on UNet2 for a test image from the OCTA-500 dataset.
-(a) Test image.
-(b) Segmentation mask predicted by UNet2.
-(c,d) Pixels selected for perturbation (in black) on the test image (we sampled $924$ (1%) and $5544$ (6%) pixels with Gray intensities above $150/255$).
-(e,f) Robust (white), non-robust (red), and unknown (black) pixels for perturbation magnitudes $e=5/255$ and dimensions $r=924\times 1$ and $r=5544\times 1$.
-RQ2: Do State-of-the-Art Deterministic Verification Techniques Scale to the Level Achieved by Our Method on Complex and High-Dimensional SSN Models?
-We applied techniques such as $\alpha$-$\beta$-CROWN (Zhou et al., 2024) and NNV (Tran et al., 2021) to our segmentation experiments using the UNet1, UNet2, and BiSeNet models in RQ1.
-However, due to the size of these models and the high perturbation levels considered, both methods encountered out-of-memory errors and failed to complete verification on the same hardware used for our approach.
-This highlights that while deterministic guarantees are often preferable, they may not always be computationally practical.
-In such cases, our probabilistic verification method for SSNs offers a scalable and effective alternative.
-RQ3: How Does the Performance of our Methodology Compare to Existing Probabilistic Verification Methods for Neural Networks in the Literature?
-In this section, we present a comparison with the works of Fiacchini and Alamo (2021); Anani et al.
-(2024) on the Cityscapes dataset Cordts et al.
-(2016).
-Since the formulation of verification guarantees differs slightly across these approaches, we first provide a general overview of the techniques proposed in Fiacchini and Alamo (2021); Anani et al.
-(2024), followed by a brief description of our own method.
+Our verification approach is designed to be scalable and efficient.
+Although it does not provide deterministic guarantees, it offers strong probabilistic assurances in regions where deterministic methods are computationally intractable.
+This effectiveness is further demonstrated through the numerical experiments included in the paper.
+Appendix A A Comparison Between application of CI and Randomized Smoothing on Robustness Analysis of SSNs.
+In this section, we present a comparison with the works of fiacchini2021probabilistic; anani2024adaptive on the Cityscapes dataset cordts2016cityscapes.
+Since the formulation of verification guarantees differs slightly across these approaches, we first provide a general overview of the techniques proposed in fiacchini2021probabilistic; anani2024adaptive, followed by a brief description of CI-based technique.
 This will clarify how the comparison can be meaningfully established.
 Overview 1.
-In Fischer et al.
-(2021), the authors introduced a probabilistic method to verify Semantic Segmentation Neural Networks.
-Inspired by Cohen et al.
-(2019), they introduced randomness via a Gaussian noise $\nu\sim\mathcal{N}(0,\sigma^{2})\in\mathbb{R}^{h\times w\times nc}$, applied to the input, and constructed a smoothed version of the segmentation model, denoted $\overline{\mathsf{SSN}}(x)(i,j)$, for each mask pixel $(i,j)$:
+In fischer2021scalable, the authors introduced a probabilistic method to verify Semantic Segmentation Neural Networks.
+Inspired by cohen2019certified, they introduced randomness via a Gaussian noise $\nu\sim\mathcal{N}(0,\sigma^{2})\in\mathbb{R}^{h\times w\times nc}$, applied to the input, and constructed a smoothed version of the segmentation model, denoted $\overline{\mathsf{SSN}}(x)(i,j)$, for each mask pixel $(i,j)$:
 $\overline{\mathsf{SSN}}(x)(i,j)=c_{A}(i,j)=\underset{c\in\mathbf{L}}{\arg\max}\ \Pr_{\nu\sim\mathcal{N}(0,\sigma^{2})}\left[\mathsf{SSN}(x+\nu)(i,j)=c\right]$
@@
 While this simplifies the formulation, it introduces the concept of abstained or uncertifiable pixels—those that cannot meet the desired confidence threshold.
-To mitigate the issue of multiple comparisons (i.e., union bounds), the authors propose statistical corrections such as the Bonferroni and Holm–Bonferroni methods CE (1936); Holm (1979).
+To mitigate the issue of multiple comparisons (i.e., union bounds), the authors propose statistical corrections such as the Bonferroni and Holm–Bonferroni methods ce1936pubblicazioni; holm1979simple.
 Assuming the set of certifiable pixels is defined as $\mathbf{CERT}=\left\{(i,j)\mid(i,j)\text{ is certifiable}\right\}$, they propose the final guarantee in the following form:
@@
 A key limitation of this technique is the presence of abstained pixels.
-The authors in Anani et al.
-(2024) addressed this by proposing an adaptive approach, $\mathsf{ADAPTIVECERTIFY}$, which reduces the number of uncertifiable pixels.
-However, the guarantees proposed in Fiacchini and Alamo (2021); Anani et al.
-(2024) apply only to the smoothed model, not the base model.
+The authors in anani2024adaptive addressed this by proposing an adaptive approach, $\mathsf{ADAPTIVECERTIFY}$, which reduces the number of uncertifiable pixels.
+However, the guarantees proposed in fiacchini2021probabilistic; anani2024adaptive apply only to the smoothed model, not the base model.
 To express this in terms of the base model, we define the following problem:
@@
 Overview 2.
-In our approach, for a given image $x$ and input set $\mathbf{B}_{r}(x)$, we perform a probabilistic reachability analysis with a $\langle\epsilon,\ell,m\rangle$ guarantee.
-Due to the presence of conservatism in our reachability technique, some mask pixels are marked as robust (i.e., certifiable), while others cover multiple classes and are considered unknown or uncertifiable.
+In CI approach, for a given image $x$ and input set $\mathbf{B}_{r}(x)$, we perform a probabilistic reachability analysis with a $\langle\epsilon,\ell,m\rangle$ guarantee.
+Due to the presence of conservatism in the reachability technique, some mask pixels are marked as robust (i.e., certifiable), while others cover multiple classes and are considered unknown or uncertifiable.
 Let $\mathbf{CONFORMAL\_CERT}=\left\{(i,j)\mid(i,j)\text{ is certifiable}\right\}$.
-Then our verification objective is:
+Then the verification objective is:
 Problem 2.
-Given an image $x$, input set $\mathbf{B}_{r}(x)$, and sampling distribution $x^{\prime}\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{B}_{r}(x)$, for coverage level $\delta_{1}=1-\epsilon\in[0,1]$, hyper-parameters $\ell,m$ and confidence level $\delta_{2}=1-\mathsf{betacdf}_{\delta_{1}}(\ell,m+1-\ell)$, we aim to show the following $\langle\epsilon,\ell,m\rangle$ guarantee, where here the set $\mathbf{CONFORMAL\_CERT}$ will be also determined through the verification process:
+Given an image $x$, input set $\mathbf{B}_{r}(x)$, and sampling distribution $x^{\prime}\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{B}_{r}(x)$, for coverage level $\delta_{1}=1-\epsilon\in[0,1]$, hyper-parameters $\ell,m$ and confidence level $\delta_{2}=1-\mathsf{betacdf}_{\delta_{1}}(\ell,m+1-\ell)$, we aim to show the following $\langle\epsilon,\ell,m\rangle$ guarantee, where the set of certifiable pixels $\mathbf{CONFORMAL\_CERT}$ will be also determined through the verification process:
 $\Pr\big[\Pr\big[\hskip-28.45274pt\bigwedge_{(i,j)\in\mathbf{CONFORMAL\_CERT}}\hskip-28.45274pt\mathsf{SSN}(x^{\prime})(i,j)=\mathsf{SSN}(x)(i,j)\hskip 8.53581pt\big]\geq\delta_{1}\big]\geq\delta_{2}$
 Comparison.
-The unknown pixels in our method are conceptually equivalent to the abstained pixels in Fiacchini and Alamo (2021); Anani et al.
-(2024).
-Thus, given the same input set $\mathbf{B}_{r}(x)$ where $r$ is provided by Fiacchini and Alamo (2021), the comparison focuses to show which technique results in fewer uncertifiable mask pixels.
-To this end, we replicate the setup of Table 1 from Anani et al.
-(2024), using the same HrNetV2 model Wang et al.
-(2020)trained on the Cityscapes dataset with the HrNetV2-W48 backbone.
-In this case study, the results of our Naive approach were approximately equivalent to those of the surrogate-based approach.
+The unknown pixels in the CI-based technique are conceptually equivalent to the abstained pixels in fiacchini2021probabilistic; anani2024adaptive.
+Thus, given the same input set $\mathbf{B}_{r}(x)$ where $r$ is provided by fiacchini2021probabilistic, the comparison focuses to show which technique results in fewer uncertifiable mask pixels.
+To this end, we replicate the setup of Table 1 from anani2024adaptive, using the same HrNetV2 model wang2020deeptrained on the Cityscapes dataset with the HrNetV2-W48 backbone.
+In this case study, the results of the Naive approach were approximately equivalent to those of the surrogate-based approach.
 Therefore, we use the Naive technique for comparison, as it is more efficient than the surrogate-based method.
-We compare the number of uncertifiable pixels produced by our naive approach against the abstained pixels reported in their results.
+We compare the number of uncertifiable pixels produced by the naive approach against the abstained pixels reported in their results.
 Our findings are summarized in Table 2, which shows the average over 200 test images.
-We also report our verification runtime for completeness.
-Table 3: Percentage of uncertifiable pixels under different $(\sigma,\tau,r)$ settings for Fiacchini and Alamo (2021), Anani et al.
-(2024), and our Naive approach.
+We also report the verification runtime for completeness.
+It is important to note that although the number of uncertifiable pixels in CI-based method is significantly smaller than in prior approaches, the CI-based formulation of guarantees also differs.
+In particular, we assume a prior distribution for the uncertainty, where $\mathcal{W}$ is considered to be uniform.
+This assumption, however, can be readily extended to worst-case distribution analysis by replacing conformal inference with robust conformal inference cauchois2020robust, a technique that strengthens guarantees at a modest cost.
+We leave this extension to future work.
+Table 2: Percentage of uncertifiable pixels under different $(\sigma,\tau,r)$ settings for fiacchini2021probabilistic, anani2024adaptive, and the Naive approach.
 The verification runtime is $210$ seconds for all experiments.
-$\sigma$ $\tau$ $r$ Fiacchini and Alamo (2021) (%) Anani et al. (2024) (%) Ours (%)
+$\sigma$ $\tau$ $r$ fiacchini2021probabilistic (%) anani2024adaptive (%) Naive approach (%)
 0.25 0.75 0.1686 7 5 0.0642
@@
 0.50 0.95 0.8224 39 28 0.0758
+Appendix B Naive Reachability Technique via Conformal Inference.
+The authors in hashemiscaling first construct a probabilistic reachable set for neural networks solely using the conformal inference, yielding a hyper-rectangular reachset.
+To achieve this, they firstly generate a calibration dataset $\mathbf{M}$ of size $m$ along with a training dataset $\mathbf{T}$ of size $t$.
+To construct the training dataset, $t$ inputs $x_{j}^{\text{train}},;j=1,2,\ldots,t$ are sampled from $\mathbf{I}$ according to any chosen distribution, denoted $x\stackrel{{\scriptstyle\mathcal{W^{\prime}}}}{{\sim}}\mathbf{I}$.
+Their corresponding outputs are then computed as $\mathbf{vec}(y_{j}^{\text{train}})=f(\mathbf{vec}(x_{j}^{\text{train}}))$.
+The resulting collection forms the training dataset, $\mathbf{T}=\left\{(x_{1}^{\text{train}},y_{1}^{\text{train}}),(x_{2}^{\text{train}},y_{2}^{\text{train}}),\ldots,(x_{t}^{\text{train}},y_{t}^{\text{train}})\right\}$.
+To construct the calibration dataset, $m$ inputs $x^{\text{calib}}_{i},;i=1,2,\ldots,m$ are sampled from $\mathbf{I}$ according to the distribution $x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}$.
+Their corresponding outputs are computed as $\mathbf{vec}(y^{\text{calib}}_{i})=f(\mathbf{vec}(x^{\text{calib}}_{i}))$.
+This yields the input/output dataset
+$\mathbf{IO}=\left\{(x_{1}^{\text{calib}},y^{\text{calib}}_{1}),(x_{2}^{\text{calib}},y^{\text{calib}}_{2}),\ldots,(x_{m}^{\text{calib}},y^{\text{calib}}_{m})\right\}.$
+This dataset is then used to compute a suitable nonconformity score $R\in\mathbb{R}_{\geq 0}$.
+Motivated by recent applications of conformal inference in time-series cleaveland2023conformal; hashemi2024statistical, the authors design the nonconformity scores to produce a hyper-rectangular set.
+For an output $\mathbf{vec}(y)=[\mathbf{vec}(y)(1),\ldots,\mathbf{vec}(y)(n)]\in\mathbb{R}^{n}$, hashemiscaling selects the calibration scores as,
+$R^{\text{calib}}_{i}=\max\left(\frac{|\mathbf{vec}(y^{\text{calib}}_{i})(1)-c(1)|}{\tau_{1}},\ldots,\frac{|\mathbf{vec}(y^{\text{calib}}_{i})(n)-c(n)|}{\tau_{n}}\right),\quad i=1,2,\ldots,m,$ (6)
+where $c=\sum_{j=1}^{t}\mathbf{vec}(y_{j}^{\text{train}})$ is the average of the outputs in $\mathbf{T}$.
+For each coordinate $k=1,2,\ldots,n$, the normalization factor $\tau_{k}$ rescales the random variables $|\mathbf{vec}(y)-c|$ and is computed from the training dataset as
+$\tau_{k}:=\max\left(\tau^{*},\ \max\big(|\mathbf{vec}(y_{1}^{\text{train}})(k)-c(k)|,\ldots,|\mathbf{vec}(y_{t}^{\text{train}})(k)-c(k)|\big)\right),$ (7)
+with $\tau^{*}=10^{-5}\left(\sum_{k=1}^{n}\sum_{j=1}^{t}|\mathbf{vec}(y_{j}^{\text{train}})(k)-c(k)|\right)/nt$ introduced to prevent division by zero.
+Finally, the calibration dataset is given by $\mathbf{M}=\{R^{\text{calib}}_{1},R^{\text{calib}}_{2},\ldots,R^{\text{calib}}_{m}\}$.
+Following the principles of conformal inference, the nonconformity scores in $\mathbf{M}$ are first sorted in ascending order.
+Without loss of generality, assume $R^{\text{calib}}_{1}<R^{\text{calib}}_{2}<\cdots<R^{\text{calib}}_{m}$.
+Now, consider a new sample $x^{\text{unseen}}$ drawn from the distribution $x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}$.
+Given the neural network architecture, the corresponding output is $\mathbf{vec}(y^{\text{unseen}})=f(\mathbf{vec}(x^{\text{unseen}}))$, which follows some distribution denoted by $y^{\text{unseen}}\sim\mathcal{Y}$.
+Through the proposed mapping for nonconformity scores, the associated random variable is
+$R^{\text{unseen}}=\max\left(\frac{|\mathbf{vec}(y^{\text{unseen}})(1)-c(1)|}{\tau_{1}},\ldots,\frac{|\mathbf{vec}(y^{\text{unseen}})(n)-c(n)|}{\tau_{n}}\right).$ (8)
+This induces another distribution, written $R\sim\mathcal{D}$.
+Both $R^{\text{unseen}}$ and the nonconformity scores $R^{\text{calib}}_{1},R^{\text{calib}}_{2},\ldots,R^{\text{calib}}_{m}$ are all i.i.d.
+samples from $\mathcal{D}$, which ensures the applicability of conformal inference.
+Consequently, for the new draw $R^{\text{unseen}}$, given a desired rank $\ell\leq m$ and miscoverage level $\epsilon$, hashemiscaling chooses the following $\langle\epsilon,\ell,m\rangle$ guarantee:
+$\Pr\left[\Pr\!\left(R^{\text{unseen}}\leq R^{\text{calib}}_{\ell}\right)>1-\epsilon\right]>1-\mathsf{betacdf}_{1-\epsilon}(\ell,m+1-\ell).$
+From Equation (8), the condition $R^{\text{unseen}}\leq R_{\ell}^{\text{calib}}$ can be expressed as the logical statement $P_{1}(\ell,m):=[R^{\text{unseen}}\leq R_{\ell}^{\text{calib}}]$.
+This condition is equivalent to requiring that every coordinate of $\mathbf{vec}(y^{\text{unseen}})$ lies within a bounded interval around $c(k)$, namely
+$P_{2}(\ell,m):=\bigwedge_{k=1}^{n}\left[c(k)-\sigma_{k}\leq\mathbf{vec}(y^{\text{unseen}})(k)\leq c(k)+\sigma_{k}\right],$ (9)
+where $\sigma_{k}=\tau_{k}R^{\text{calib}}_{\ell}$.
+This describes a hyper-rectangular region that serves as the reachset for unseen outputs of the neural network sampled from $x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}$.
+Since $P_{1}(\ell,m)$ carries the $\langle\epsilon,\ell,m\rangle$ guarantee, the equivalent hyper-rectangular reachset inherits the same coverage guarantee for unseen outputs.
+Advantages and Disadvantages for Application of CI on SSN
+Reasoning directly about the output distribution $y\sim\mathcal{Y}$ resulting from inputs $x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}$ is generally infeasible due to the strong nonlinearity of neural networks.
+Consequently, providing probabilistic coverage guarantees over the network’s output space becomes highly challenging.
+Conformal inference (CI) offers a key advantage in this setting: its guarantee is distribution free, meaning they remain valid for any distribution family that adequately represents the calibration dataset.
+This property makes CI particularly suitable for neural network reachability analysis, where output distributions are often complex.
+Despite this, applying CI to high-dimensional outputs introduces significant difficulties.
+The scalar nonconformity scores $R$ in the calibration dataset are defined based on the network outputs $y\sim\mathcal{Y}$, but in high dimensions, the family of distributions $y\sim\mathcal{Y}$ that can generate the calibration distribution $R\sim\mathcal{D}$ grows dramatically in size.
+As a result, CI is robust to all the members of this family and thus tends to produce overly conservative guarantees, which limits its practical applicability to networks like semantic segmentation networks (SSNs) that generate extremely high-dimensional outputs.
+To overcome this limitation, hashemiscaling propose a new learning-based reachability analysis method for SSNs scalable for high dimensional settings.
+Appendix C Toy Example
+In this section we provide a comparison against the naïve approach from hashemiscaling. Specifically, we consider the toy example proposed in hashemiscaling provided below.
+Example 1
+Consider a feedforward $\mathrm{ReLU}$ neural network with $60$ hidden-layers of width $100$, an input layer of dimension $784$ and an output layer of dimension $2$.
+The set of inputs is $\mathbf{I}:=[0,1]^{784}$ and we generate both calibration and train datasets by sampling $\mathbf{I}$ uniformly.
+e.g., $\mathcal{W},\mathcal{W}^{\prime}$ are both uniform distributions.
+We plan to generate reachable sets for this model that satisfies the following $\langle\epsilon,\ell,m\rangle$ guarantee with our technique and naive approach.
+$x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}\ \ \Rightarrow\ \ \Pr\left[\ \ \Pr[\ \ f(x)\in\mathbf{S}_{1}\ \ ]>0.9999\ \ \right]>0.9999995.$ (10)
+To solve this problem hashemiscaling generates a calibration dataset of size $m=200,000$ and a train dataset of size $t=10,000$, they also consider the rank $\ell=199,998$ and target the miscoverage level of $\epsilon=0.0001$.
+In this case they compute the reachable set, $\mathbf{S}_{1}$ presented in Figure 3 in orange.
+On the other hand, we collect a calibration dataset of size m=200,000m=200{,}000 and compare the resulting bounds.
+For this comparison, we employ a clipping block that projects with $l=\infty$.
+We constructed the convex hull using a training dataset of size $t=4000$.
+In addition, we sampled a separate dataset of size $t^{\prime}=10000$ to compute the normalization factors (see Eq.
+(7)) prior to performing conformal inference for obtaining the inflating hyper-rectangle.
+The resulting reachable set $\mathbf{S}_{3}$, together with the outcome of the naïve approach, in the presence of $10^{6}$ new simulations, $f(x),x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}$, are shown in Figure 3.
+Our calculations show that the proportion of points lying outside of $\mathbf{S}_{1}$ and $\mathbf{S}_{3}$ are 0.000013 and 0.000012, respectively, which aligns well with the proposed $\langle\epsilon,\ell,m\rangle$ guarantees.
+Figure 3: The figure compares two reachable sets with the same $\langle\epsilon,\ell,m\rangle$ guarantees.
+$S_{1}$ is obtained using the naïve approach, while $S_{3}$ is produced by the surrogate-based technique, illustrating the accuracy of the latter.
+Appendix D Additional Research Questions
+Table 3: This table reports the bound ratio, $\mathsf{bound\_ratio}$, and the empirical miscoverage level of our reachset, $\hat{\epsilon}$ for a selection of test images, perturbation level and perturbation dimensions.
+We performed $10^{6}$ inferences to estimate the conservatism.
+Dataset Model Image name # Perturbed Pixels $e$ $\mathsf{bound\_ratio}$ $\hat{\epsilon}$
+Camvid BiSeNet 0001TP_008790.png $41471\times 3$ $5/255$ 0.5328 3.08e-6
+Figure 4: Shows the average run time versus perturbation level (top row) and perturbation dimension $r$ (bottom row).
+In the top row, the image is entirely perturbed within an $\ell_{2}$ ball for CamVid and Lung Segmentation and $\ell_{\infty}$ ball for OCTA-500.
+In the bottom row, we have darkening adversary where $e$ is fixed at $5/255$ across all experiments.
+The runtimes are averaged over 200 test images from datasets.
+(a) Image, 0001TP_008790.png from test dataset
+(b) The Mask prediction using the BiSeNet model
+(c) Image perturbed over $6912$ darkened pixels.
+(d) Image perturbed over $41472$ darkened pixels
+(e) Unknown and non-robust masks for darkening adversary with $e=5/255$ and $r=6192\times 3$.
+(f) Unknown and non-robust masks for darkening adversary with $e=5/255$ and $r=41472\times 3$.
+Figure 5: Visualization for verification on BiSeNet for a test image from the CamVid dataset.
+(a) The test image used for verification.
+(b) The segmentation mask predicted by BiSeNet for this image.
+(c,d) The pixels selected for perturbation (in black) on the test image (We sampled $6192$ (1% of) and $41472$ (6% of) pixels where the R, G, and B intensities were all above $150/255$, forming a perturbation set $\mathbf{I}\subset\mathbb{R}^{6192\times 3}$ and $\mathbf{I}\subset\mathbb{R}^{41472\times 3}$ respectively).
+(e,f) Display the robust (white), non-robust (red), and unknown (black) pixels for the perturbation set $\mathbf{I}$ as described in Figure 1, with perturbation magnitudes $e=5/255$ and perturbation dimension $r=6192\times 3$ and $r=41472\times 3$, respectively.
+(a) Image, CHNCXR_0005_0.png from test dataset
+(b) The Mask prediction generated via the model UNet1
+(c) Image perturbed over $2621$ darkened pixels.
+(d) Image perturbed over $15726$ darkened pixels.
+(e) Unknown and non-robust masks for darkening adversary with $e=5/255$ and $r=2621\times 1$.
+(f) Unknown and non-robust masks for darkening adversary with $e=5/255$ and $r=15726\times 1$.
+Figure 6: Visualization for verification on UNet1 for a test image from the Lung Segmentation dataset.
+(a) The test image used for verification.
+(b) The segmentation mask predicted by UNet1 for this image.
+(c,d) The pixels selected for perturbation (in black) on the test image (We sampled 26212621(1% of) and $15726$ (6% of) pixels where the Gray intensities were above $150/255$, forming a perturbation set $\mathbf{I}\subset\mathbb{R}^{2621\times 1}$ and $\mathbf{I}\subset\mathbb{R}^{15726\times 1}$ respectively).
+(e,f) Display the robust (white), non-robust (red), and unknown (black) pixels for the perturbation set $\mathbf{I}$ as described in Figure 1, with perturbation magnitudes $e=5/255$ and perturbation dimnesion $r=2621\times 1$ and $r=15726\times 1$, respectively.
+(a) Image, 10491.bmp from OCTA-500 dataset.
+(b) The Mask prediction generated by the model UNet2
+(c) Image perturbed over $924$ darkened pixels.
+(d) Image perturbed over $5544$ darkened pixels.
+(e) Unknown and non-robust masks for darkening adversary with $e=5/255$ and $r=924\times 1$.
+(f) Unknown and non-robust masks for darkening adversary with $e=5/255$ and $r=5544\times 1$.
+Figure 7: Visualization for verification on UNet2 for a test image from the OCTA-500 dataset.
+(a) The test image used for verification.
+(b) The segmentation mask predicted by UNet2 for this image.
+(c,d) The pixels selected for perturbation (in black) on the test image (We sampled 924924(1% of) and $5544$ (6% of) pixels where the Gray intensities were above $150/255$, forming a perturbation set $\mathbf{I}\subset\mathbb{R}^{924\times 1}$ and $\mathbf{I}\subset\mathbb{R}^{5544\times 1}$ respectively).
+(e,f) Display the robust (white), non-robust (red), and unknown (black) pixels for the perturbation set $\mathbf{I}$ as described in Figure 1, with perturbation magnitudes $e=5/255$ and perturbation dimension $r=924\times 1$ and $r=5544\times 1$, respectively.
+RQ2: Do State-of-the-Art Deterministic Verification Techniques Scale to the Level Achieved by Our Method on Complex and High-Dimensional SSN Models?
+We applied techniques such as $\alpha$-$\beta$-CROWN (zhou2024scalable) and NNV (tran2021robustness) to our segmentation experiments using the UNet1, UNet2, and BiSeNet models in RQ1.
+However, due to the size of these models and the high perturbation levels considered, both methods encountered out-of-memory errors and failed to complete verification on the same hardware used for our approach.
+This highlights that while deterministic guarantees are often preferable, they may not always be computationally practical.
+In such cases, our probabilistic verification method for SSNs offers a scalable and effective alternative.